### v1

Abstract
This paper investigates the continuous-time limit of score-driven models with long memory. By extending score-driven models to incorporate infinite-lag structures with coefficients exhibiting heavy-tailed decay, we establish their weak convergence, under appropriate scaling, to fractional Ornstein-Uhlenbeck processes with Hurst parameter $H<1/2$. When score-driven models are used to characterize the dynamics of volatility, they serve as discrete-time approximations for rough volatility. We present several examples, including EGARCH($\infty$) whose limits give rise to a new class of rough volatility models. Building on this framework, we carry out numerical simulations and option pricing analyses, offering new tools for rough volatility modeling and simulation.
Keywords: Score-driven models, Rough volatility, Weak convergence, Option pricing
2020 MSC: 60F17, 60G22, 62M10, 91B70
1 Introduction
Score-driven models, also known as generalized autoregressive score models, were proposed by Creal et al., (2013) and Harvey, (2013) as a class of observation-driven time series models in which parameter updates are driven by the score—the gradient of the log-likelihood function. By directly linking the score of the data distribution to parameter dynamics in a unified framework, these models inherit flexibility and generality, encompassing many classical models such as GARCH (Bollerslev, , 1986) and ACD (Engle and Russell, , 1998) models. Theoretically, score-driven models have been proven optimal in minimizing the Kullback-Leibler divergence between the true distribution and the postulated distribution (Gorgi et al., , 2024). Moreover, in statistical inference, the model’s closed-form likelihood function facilitates efficient maximum likelihood estimation, exhibiting lower computational costs compared to parameter-driven models. In application, they also demonstrate robust empirical performance across diverse fields such as economics, finance and biology. Within just a few years, the literature on SD models has grown to nearly 400 and can be found online at http://www.gasmodel.com/gaspapers.htm.
As a remarkably influential model in time series analysis, it naturally raises the question: what would its continuous-time counterpart look like? Buccheri et al., (2021) have investigated its continuous-time limit, obtaining a two-dimensional diffusion process. Similarly, Wu and He, (2024) investigated the continuous-time limit of the quasi score-driven volatility model proposed by Blasques et al., (2023), and derived a correlated stochastic volatility model. However, the score-driven models they studied only incorporate first-order or finite-order lags—these are all Markov processes. In fact, when the model was first introduced in Creal et al., (2013), the question was raised as to whether it could be extended to an infinite-lag form. Janus et al., (2014) and Opschoor and Lucas, (2019) pursued such extensions by proposing long memory score-driven models. This naturally leads to the question: What would be the continuous-time counterpart of such a long memory version?
On the other hand, since Gatheral et al., (2018) discovered that volatility is rough, rough volatility models have gained popularity in financial modeling. These models are characterized by the use of fractional Brownian motion with a Hurst parameter $H<1/2$, whose paths are rougher than those of Brownian motion—hence the name. El Euch et al., (2018) investigated the microstructure of rough volatility by employing Hawkes processes to model the arrival dynamics in limit order books. This theoretical basis originates from the discovery of Jaisson and Rosenbaum, (2016) that the scaling limit of nearly unstable heavy-tailed Hawkes processes is a rough Cox-Ingersoll-Ross (CIR) process. It is worth noting that Hawkes processes are self-exciting (similar to observation-driven mechanisms) and exhibit long memory, which provides inspiration for us. In particular, Cai et al., (2024) studied the discrete-time counterpart of the Hawkes process—the INAR($\infty$) process—and obtained similar convergence results. This motivates our investigation into the scaling limit of score-driven models when extended to incorporate long memory.
This paper establishes that a sequence of long memory score-driven models, under appropriate scaling, converges weakly to a rough Ornstein-Uhlenbeck (OU) process. By modeling time-varying parameters as log-volatility (as in EGARCH-type models), we obtain the limit a class of rough volatility models. Consequently, long memory score-driven volatility models can serve as discrete-time approximations for such rough volatility models, facilitating efficient Monte Carlo simulations for financial applications. To the best of our knowledge, apart from Wang and Cui, (2025)’s approach using INAR($\infty$) processes to approximate the rough Heston model, this study presents one of the few methodologies employing time series models to approximate rough volatility processes. Moreover, we present specific examples of long memory score-driven volatility models to further demonstrate the practical versatility and broad applicability of our theoretical results.
The remainder of the paper is structured as follows. Section 2 introduces the long memory score-driven models that we focus on and provides a heuristic derivation of their continuous-time limits. In Section 3, we establish the main convergence results and presents their proofs. Section 4 applies these results to volatility modeling, illustrating with the Gamma-GED-EGARCH($\infty$) model and its extension. We conducts corresponding approximations of rough volatility models and option pricing. Section 5 concludes our findings and provides an outlook for future research.
2 Long memory score-driven models
Since Engle, (1982) discovered that economic data exhibit conditional heteroskedasticity, it has gradually been recognized that certain characteristics of time series are not immutable. Specifically, let $\{y_{n}\}_{n\in\mathbb{N}}$ be a time series and $\mathcal{F}{n}=\sigma(y_{n},y_{n-1},\ldots,y_{0})$ be the $\sigma$-algebra it generates. The conditional density
$y_{n}|\mathcal{F}_{n-1}\sim p(y_{n}|\lambda_{n},\mathcal{F}_{n-1}),$
with a time-varying parameter $\lambda_{n}$. To address this issue, Cox et al., (1981) proposed two modeling approaches: observation-driven models, where the dynamics of the time-varying parameters are governed by past observations of the series itself, and parameter-driven models, in which the parameter dynamics are driven by another stochastic process.
The score-driven model is a class of observation-driven models whose key idea is to update time-varying parameters based on the score of the observation. Specifically, the parameter update follows the form:
$\lambda_{n+1}=\omega+\beta\lambda_{n}+\alpha S(\lambda_{n})\nabla_{n}.$ (2.1)
where $\nabla_{n}=\frac{\partial\log p(y_{n}|\lambda_{n},\mathcal{F}{n-1})}{\partial\lambda_{n}}$ denotes the partial derivative of the log-density function of observations with respect to the parameter, commonly referred to as the (Fisher) score in statistics. Consequently, these models are aptly termed score-driven models. The function $S(\cdot)$ is known as the scaling function, which accounts for the curvature of the log-density function. It is typically chosen as the negative exponent of the conditional Fisher information, i.e., $S(\lambda_{n})=[\mathbb{E}(\nabla_{n}^{2}|\mathcal{F}{n-1})]^{-a}$, with common choices for $a$ being 0, $1/2$, or $1$. To illustrate, when the density function $p(\cdot)$ is chosen to be that of a normal distribution and the parameter $\lambda$ is identified as the conditional variance $\sigma^{2}$, with $a=1$, the score-driven model reduces to the classical GARCH model.
The equation given by (2.1) represents a first-order score-driven model, as the relationship between the score and the parameter is lagged by only one period. This framework can be extended to higher orders, such as $p$-th order or even infinite order, as in Janus et al., (2014),
$\lambda_{n}=\omega+\sum_{i=1}^{n}\phi_{i}S(\lambda_{n-i})\nabla_{n-i}.$
Noting that our time index starts from 0, the summation at time $n$ includes only $n$ terms. The term “infinite” refers to the dependence on the entire history. Additionally, we omit the autoregressive term, as it can be absorbed into the coefficients $\phi_{i}$ through the moving average representation. This formulation appears similar to the ARCH($\infty$) model, but the key difference lies in the fact that the infinite-order lagged term $S\nabla$ in the score-driven model is a martingale difference, whereas the $y^{2}$ term in the ARCH($\infty$) model is not. In other words, rewriting the ARCH($\infty$) model in a form driven by martingale differences gives
$\sigma^{2}_{n}=\omega+\sum_{i=1}^{n}\phi_{i}y^{2}_{n-i}=\omega+\sum_{i=1}^{n}\phi_{i}\sigma^{2}_{n-i}+\sum_{i=1}^{n}\phi_{i}(y^{2}_{n-i}-\sigma^{2}_{n-i}),$
which includes an additional infinite autoregressive component. To address this, Opschoor and Lucas, (2019) proposed the following formulation:
$\lambda_{n}=\omega+\sum_{i=1}^{n}\phi_{i}[\lambda_{n-i}+S(\lambda_{n-i})\nabla_{n-i}].$ (2.2)
The long memory score-driven model we considered henceforth refers to (2.2). To investigate its scaling limit, we define a sequence of processes:
$\begin{split}&y_{t}^{(n)}|\mathcal{F}_{t-1}^{(n)}\sim p\left(y_{t}^{(n)}|\lambda_{t}^{(n)},\mathcal{F}^{(n)}_{t-1}\right),\\ &\lambda_{t}^{(n)}=\omega_{n}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+S(\lambda_{t-i}^{(n)})\nabla_{t-i}^{(n)}\right],\end{split}$ (2.3)
where $\phi_{i}^{(n)}=a_{n}\phi_{i}>0$. Let $a_{n}>0$ and $\uparrow 1$ as $n\to\infty$, $\|\phi\|=\sum_{i=1}^{\infty}\phi_{i}=1$, so that $\|\phi^{(n)}\|\uparrow 1$. This case is referred to as nearly unstable. Specifically, we set $\phi_{i}\sim K/i^{1+\alpha}$, exhibiting power-law decay, where $\alpha\in(1/2,1)$, a form referred to as “heavy-tailed” in Jaisson and Rosenbaum, (2016). This condition is critical for the limit to generate roughness. The sigma-algebra is defined as $\mathcal{F}^{(n)}_{t}=\sigma(\lambda_{0}^{(n)},y_{i}^{(n)},\lambda_{i+1}^{(n)},i=0,1,\ldots,t)$, since $\lambda_{t+1}^{(n)}$ is determined when $\{y_{i}^{(n)}\}_{i\leq t}$ are known.
Our primary focus is on the dynamics of time-varying parameter $\lambda_{t}^{(n)}$, and we aim to study its scaling limiting behavior, which we expect to converge to a SDE driven by a rough fractional Brownian motion. To build intuition for this convergence, we first rewrite the equation. For notation simplicity, the superscript $n$ is sometimes omitted without ambiguity, and the terms $a_{n}$, $\omega_{n}$ serving as a reminder that we are dealing with a sequence indexed by $n$. Define $M_{t}=\sum_{i=0}^{t}S(\lambda_{i})\nabla_{i}$, with $M_{-1}=0$. Clearly, $M$ is a martingale, since
$\mathbb{E}(M_{t}-M_{t-1}|\mathcal{F}_{t-1})=\mathbb{E}[S(\lambda_{t})\nabla_{t}|\mathcal{F}_{t-1}]=S(\lambda_{t})\mathbb{E}[\nabla_{t}|\mathcal{F}_{t-1}]=0.$
Then,
$\begin{split}\lambda_{t}&=\omega_{n}+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}[\lambda_{i}+S(\lambda_{i})\nabla_{i}]\\ &=\omega_{n}+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}(M_{i}-M_{i-1})+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}\lambda_{i}\\ &=\omega_{n}+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}(M_{i}-M_{i-1})+\sum_{i=0}^{t-1}\psi_{t-i}^{(n)}\left[\omega_{n}+\sum_{j=0}^{i-1}a_{n}\phi_{i-j}(M_{j}-M_{j-1})\right].\end{split}$
The final equality follows from the transformation between AR($\infty$) and MA($\infty$), that is,
$\lambda_{t}=\varepsilon_{t}+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}\lambda_{i}\quad\Longleftrightarrow\quad\lambda_{t}=\varepsilon_{t}+\sum_{i=0}^{t-1}\psi^{(n)}_{t-i}\varepsilon_{i}.$
Here, $\psi^{(n)}=\sum_{k=1}^{\infty}(a_{n}\phi)^{*k}$ represents the sum of the $k$-fold convolution of $\phi^{(n)}$. By rearranging the expression and exchanging the order of the double summation, we obtain
$\lambda_{t}=\omega_{n}+\sum_{i=0}^{t-1}\psi_{t-i}^{(n)}\omega_{n}+\sum_{i=0}^{t-1}a_{n}\phi_{t-i}(M_{i}-M_{i-1})+\sum_{j=0}^{t-2}(M_{j}-M_{j-1})\sum_{i=j+1}^{t-1}a_{n}\phi_{i-j}\psi_{t-i}^{(n)}.$
Note that
$\sum_{i=j+1}^{t-1}a_{n}\phi_{i-j}\psi_{t-i}^{(n)}=(\phi^{(n)}*\psi^{(n)})_{t-j},$
and
$\phi^{(n)}*\psi^{(n)}=\sum_{k=1}^{\infty}(\phi^{(n)})^{*{k+1}}=\sum_{k=2}^{\infty}(\phi^{(n)})^{*{k}}=\psi^{(n)}-\phi^{(n)},$
we have
$\begin{split}\lambda_{t}&=\omega_{n}+\sum_{i=0}^{t-1}\psi_{t-i}^{(n)}\omega_{n}+a_{n}\phi_{1}(M_{t-1}-M_{t-2})+\sum_{j=0}^{t-2}\psi_{t-j}^{(n)}(M_{j}-M_{j-1})\\ &=\omega_{n}+\sum_{i=0}^{t-1}\psi_{t-i}^{(n)}\omega_{n}+\sum_{j=0}^{t-1}\psi_{t-j}^{(n)}(M_{j}-M_{j-1}).\end{split}$ (2.4)
Since $\mathbb{E}[\lambda_{t}^{(n)}]\to\omega_{n}/(1-a_{n})$ as $t\to\infty$, we apply a rescaling to $\lambda_{t}^{(n)}$ and define a new process
$\Lambda_{t}^{(n)}:=\frac{(1-a_{n})\theta}{\omega_{n}}\lambda_{\lfloor nt\rfloor}^{(n)},\quad t\in[0,1],$
where $\theta>0$ is a constant controlling the mean level of $\Lambda_{t}^{(n)}$. We aim to investigate the limiting behavior of $\Lambda_{t}^{(n)}$ as $n\to\infty$.
Let $\Delta M_{k}=M_{k}-M_{k-1}$. Assuming
$\mbox{Var}[\Delta M_{k}|\mathcal{F}_{k-1}]=S^{2}(\lambda_{k})\mathbb{E}[(\nabla_{k})^{2}|\mathcal{F}_{k-1}]=U^{2}(\lambda_{k}).$
It follows from (2.4) that
$\begin{split}\Lambda_{t}^{(n)}&=(1-a_{n})\theta+(1-a_{n})\theta\sum_{i=0}^{\lfloor nt\rfloor-1}\psi_{\lfloor nt\rfloor-i}^{(n)}+\frac{1-a_{n}}{\omega_{n}}\theta\sum_{j=0}^{\lfloor nt\rfloor-1}\psi_{\lfloor nt\rfloor-j}^{(n)}\Delta M_{j}^{(n)}\\ &=(1-a_{n})\theta+\theta\sum_{i=1}^{\lfloor nt\rfloor}(1-a_{n})\psi_{i}^{(n)}+\frac{\theta}{\sqrt{n}\omega_{n}}\sum_{j=0}^{\lfloor nt\rfloor-1}n(1-a_{n})\psi_{\lfloor nt\rfloor-j}^{(n)}U(\lambda_{j}^{(n)})\frac{\Delta M_{j}^{(n)}}{U(\lambda_{j}^{(n)})\sqrt{n}}.\end{split}$ (2.5)
Conceivably, when $n\to\infty$,
$\sum_{j=0}^{\lfloor nt\rfloor-1}\frac{\Delta M_{j}^{(n)}}{U(\lambda_{j}^{(n)})\sqrt{n}}\Rightarrow W_{t},$
where $\{W_{t}\}_{t\geq 0}$ is a standard Brownian motion. And according to Cai et al., (2024), in that case $1/2<\alpha<1$, when $1-a_{n}\sim n^{-\alpha}$ tends to zero and the coefficients $\phi_{i}\sim K/i^{1+\alpha}$ exhibit power-law decay, the following weak convergence holds:
$n(1-a_{n})\psi^{(n)}_{\lfloor nx\rfloor}\to f^{\alpha,\kappa}(x):=\kappa x^{\alpha-1}E_{\alpha,\alpha}(-\kappa x^{\alpha}),\quad\mbox{as}\ n\to\infty,$ (2.6)
where $\kappa>0$ is a constant, and
$E_{\alpha,\beta}(x)=\sum_{n=0}^{\infty}\frac{x^{n}}{\Gamma(\alpha n+\beta)}$
is the Mittag-Leffler function. A property of the function is that it decays to zero as $x\to$-$\infty$, and $E_{\alpha,\alpha}(0)=\frac{1}{\Gamma(\alpha)}$. Thus, in terms of singularity, the behavior of $f^{\alpha,\kappa}(t-s)$ is analogous to that of the Riemann-Liouville kernel $(t-s)^{\alpha-1}$.
Therefore, informally and intuitively, as $n\to\infty$, the limit of (2.5) satisfies the following SDE:
$\Lambda_{t}=\theta\int_{0}^{t}f^{\alpha,\kappa}(s){\rm d}s+\nu\int_{0}^{t}f^{\alpha,\kappa}(t-s)U(\Lambda_{s}){\rm d}W_{s}.$ (2.7)
In fact, similar to Proposition 4.10 in El Euch et al., (2018), the SDE (2.7) has the same solution as the SDE
$\Lambda_{t}=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\kappa\nu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}U(\Lambda_{s}){\rm d}W_{s}.$ (2.8)
Since $\alpha-1\in(-1/2,0)$, the term $\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}$ corresponds to a fractional Brownian motion with Hurst parameter $H<1/2$, which means (2.8) falls into the rough case. It is worth mentioning that when $U(x)=\sqrt{x}$, this actually corresponds to the rough fractional CIR process described in Jaisson and Rosenbaum, (2016).
3 Main results
Following the heuristic derivation above, in this section, we state our main theorem and present its proof. We begin with the following assumptions.
Assumption 1.
There exists a constant $\gamma<\infty$, such that $U(x)\to\gamma$, as $x\to\infty$.
Assumption 2.
The sequences $a_{n}$ and $\omega_{n}$ have the following asymptotic behavior as $n\to\infty$:
$1-a_{n}\sim n^{-\alpha},\quad\omega_{n}\sim n^{-\frac{1}{2}}.$
Assumption 3.
For each $n$, the martingale difference sequence $\{\Delta M_{k}^{(n)}\}_{k\in\mathbb{N}}$ is strictly stationary. Furthermore, for any $p\geq 2$,
$\sup_{n\in\mathbb{N}}\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p}<\infty.$
Then, we have the following theorem.
Theorem 1.
Under Assumptions 1-3, the rescaled long memory score-driven parameter process
$\Lambda_{t}^{(n)}:=\frac{1-a_{n}}{\omega_{n}}\theta\lambda^{(n)}_{\lfloor nt\rfloor},\quad t\in[0,1]$
converges weakly in the Skorohod space $\mathcal{D}([0,1])$ to a rough fractional diffusion process given by the following SDE:
$\Lambda_{t}=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\gamma\kappa\nu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}.$ (3.1)
Where $\kappa=\lim_{n\to\infty}\frac{\alpha n^{\alpha}(1-a_{n})}{K\Gamma(1-\alpha)}$ and $\nu=\lim_{n\to\infty}\frac{\theta}{\sqrt{n}\omega_{n}}$. The existence of these limits is guaranteed by Assumption 2.
In fact, we restrict our attention to the case of rough OU process in (2.8). This is because our proof strategy fundamentally differs from that of Jaisson and Rosenbaum, (2016) or Wang and Cui, (2025): while they show that the Hawkes process (equivalent to the integrated intensity in probability) converges to a integrated rough CIR process, we directly establish the convergence of the parameter process itself to a rough process, rather than its integral. The difficulty in dealing with the process itself lies in verifying its tightness and we attempt to circumvent this by applying the continuous mapping theorem.
To this end, we regard the rough diffusion process as the image of a diffusion process under the generalized fractional operator (GFO), which is inspired by Horvath et al., (2024). Specifically, for $\lambda\in(0,1)$ and $\alpha\in(-\lambda,1-\lambda)$, $\mathcal{G}^{\alpha}$ is the GFO associated to kernel
$g\in\mathcal{L}^{\alpha}:=\left\{g\in C^{2}((0,1]):\left|\frac{g(u)}{u^{\alpha}}\right|,\left|\frac{g^{\prime}(u)}{u^{\alpha-1}}\right|\ \text{and}\ \left|\frac{g^{\prime\prime}(u)}{u^{\alpha-2}}\right|\ \mbox{are bounded}\right\}$
defined on $\lambda$-Hölder space $\mathcal{C}^{\lambda}([0,1])$,
$(\mathcal{G}^{\alpha}f)(t)=\left\{\begin{aligned} &\int_{0}^{t}(f(s)-f(0))\frac{{\rm d}}{{\rm d}t}g(t-s){\rm d}s,\quad\text{if}\ \alpha\geq 0,\\ &\frac{{\rm d}}{{\rm d}t}\int_{0}^{t}(f(s)-f(0))g(t-s){\rm d}s,\quad\text{if}\ \alpha<0.\end{aligned}\right$.
When $g(x)=x^{\alpha}$, this reduces to the classical Riemann-Liouville operator. In this paper, we choose $g(x)=f^{\alpha,\kappa}(x)\in\mathcal{L}^{\alpha-1}$ and focus on the case $\alpha\in(1/2,1)$. Consequently, the corresponding GFO actually becomes
$(\mathcal{G}^{\alpha-1}f)(t)=\frac{{\rm d}}{{\rm d}t}\int_{0}^{t}(f(s)-f(0))f^{\alpha,\kappa}(t-s){\rm d}s.$ (3.2)
Define $Y_{t}:=\theta t+\nu\gamma W_{t}$, and
$\Lambda_{t}=\theta\int_{0}^{t}f^{\alpha,\kappa}(s){\rm d}s+\nu\gamma\int_{0}^{t}f^{\alpha,\kappa}(t-s){\rm d}W_{s}$
is actually has the same solution as the (3.1). Consider the partition $\mathcal{T}:=\{t_{i}=i/n,i=0,1,\ldots,n\}$. We then define
$Y_{t}^{(n)}=\theta t+\frac{\theta}{n\omega_{n}}\left[(1-nt+\lfloor nt\rfloor)M_{\lfloor nt\rfloor-1}^{(n)}+(nt-\lfloor nt\rfloor)M_{\lfloor nt\rfloor}^{(n)}\right],$ (3.3)
$\widetilde{\Lambda}_{t}^{(n)}=\sum_{i=0}^{\lfloor nt\rfloor-1}\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{i}^{(n)}\right)\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s+\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{\lfloor nt\rfloor}^{(n)}\right)\int_{t_{\lfloor nt\rfloor}}^{t}f^{\alpha,\kappa}(t-s){\rm d}s.$ (3.4)
The subsequent proof will proceed in four steps, as illustrated by the following diagram, which provides a clear path.
  1. (i)
    Establishing the representation $\Lambda_{t}=\mathcal{G}^{\alpha-1}Y_{t}$, and $\widetilde{\Lambda}_{t}^{(n)}=\mathcal{G}^{\alpha-1}Y_{t}^{(n)}$;
  2. (ii)
    Proving the weak convergence $Y^{(n)}\Rightarrow Y$ in Hölder space $\mathcal{C}^{\lambda}$ for $\lambda<1/2$;
  3. (iii)
    Demonstrating that $\mathcal{G}^{\alpha-1}$ is a continuous operator from $\mathcal{C}^{\lambda}$ to $\mathcal{C}^{\lambda+\alpha-1}$, which yields the convergence $\widetilde{\Lambda}_{t}^{(n)}\Rightarrow\Lambda_{t}$ in $\mathcal{C}^{\lambda+\alpha-1}$, thus in $\mathcal{D}$;
  4. (iv)
    Finally, by proving $\widetilde{\Lambda}_{t}^{(n)}$ and $\Lambda_{t}^{(n)}$ are asymptotically indistinguishable, we obtain $\Lambda_{t}^{(n)}\Rightarrow\Lambda_{t}$.
$Y^{(n)}_{t}$$Y_{t}$$(\mathcal{C}^{\lambda}([0,1]),\|\cdot\|_{\lambda})$$\Lambda^{(n)}_{t}$$\widetilde{\Lambda}^{(n)}_{t}$$\Lambda_{t}$$(\mathcal{C}^{\lambda+\alpha-1}([0,1]),\|\cdot\|_{\lambda+\alpha-1})$$\scriptstyle{\|\cdot\|_{\lambda}}$$\scriptstyle{\mathcal{G}^{\alpha-1}}$$\scriptstyle{\mathcal{G}^{\alpha-1}}$$\scriptstyle{\mathcal{G}^{\alpha-1}}\scriptstyle{\approx}$$\scriptstyle{\|\cdot\|_{\lambda+\alpha-1}}$
Before proceeding with the main proof, we first recall the key convergence result (2.6) for the Mittag-Leffler function, and extend this result to $L^{2}$-convergence for our subsequent arguments.
3.1 Converges to $f^{\alpha,\kappa}(x)$
Jaisson and Rosenbaum, (2016) prove that the mean of geometric sums converges in distribution to a random variable with density function $f^{\alpha,\kappa}(x)$. Our setting differs slightly because the summed terms follow a discrete distribution.
Specifically, we define the sequence of random variables
$G_{n}=\frac{\sum_{k=1}^{I_{n}}U_{k}}{n},$
where $I_{n}$ follows a geometric distribution with parameter $1-a_{n}$, that is, $\mathbb{P}(I_{n}=i)=a_{n}^{i-1}(1-a_{n})$; $U_{k}$} is a sequence of i.i.d. random variables with distribution $\phi$, that is, $\mathbb{P}(U_{1}=i)=\phi_{i}$. Then, the random variable $G_{n}$, called the mean of geometric sums, has its law given by
$\begin{split}\mathbb{P}(G_{n}=i)&=\sum_{j=1}^{\infty}\mathbb{P}(\sum_{k=1}^{j}U_{k}=ni)\cdot\mathbb{P}(I_{n}=j)=(1-a_{n})\sum_{j=1}^{\infty}\phi^{*j}_{ni}a_{n}^{j-1}\\ &=\frac{1-a_{n}}{a_{n}}\sum_{j=1}^{\infty}(a_{n}\phi)^{*j}_{ni}=\frac{(1-a_{n})\psi^{(n)}_{ni}}{a_{n}},\quad\text{for}\ ni\in\mathbb{N}^{+}.\end{split}$
Thus, its distribution function is $F_{n}(x)=\sum_{i=1}^{\lfloor nx\rfloor}\frac{(1-a_{n})\psi^{(n)}_{i}}{a_{n}}$, density function $\rho_{n}(x)=\frac{n(1-a_{n})}{a_{n}}\psi^{(n)}_{\lfloor nx\rfloor}$. Considering the characteristic function of $G_{n}$.
$\begin{split}\hat{\rho}_{n}(u)&=\mathbb{E}({\rm e}^{-{\rm i}uG_{n}})=\mathbb{E}({\rm e}^{-\frac{{\rm i}u}{n}\sum_{k=1}^{I_{n}}U_{k}})=\sum_{j=1}^{\infty}\mathbb{E}({\rm e}^{-\frac{{\rm i}u}{n}\sum_{k=1}^{j}U_{k}})\mathbb{P}(I_{n}=j)\\ &=(1-a_{n})\sum_{j=1}^{\infty}[\mathbb{E}({\rm e}^{-\frac{{\rm i}u}{n}U_{k}})]^{j}a_{n}^{j-1}=\frac{1-a_{n}}{a_{n}}\sum_{j=1}^{\infty}[a_{n}\hat{\phi}(u/n)]^{j}\\ &=\frac{1-a_{n}}{a_{n}}\frac{a_{n}\hat{\phi}(u/n)}{1-a_{n}\hat{\phi}(u/n)}=\frac{\hat{\phi}(u/n)}{1-\frac{a_{n}}{1-a_{n}}[\hat{\phi}(u/n)-1]},\end{split}$
where $\hat{\phi}$ denotes the Fourier transform of $\phi$, or equivalently, the characteristic function of the random variable $U_{1}$. Since $\phi_{i}\sim\frac{K}{i^{1+\alpha}}$, by the Karamata-Tauberian theorem (see Cai et al., (2024)), we can characterize the behavior of $\hat{\phi}$ near the origin,
$1-\hat{\phi}(u)\underset{u\to 0}{\sim}\frac{K\Gamma(1-\alpha)}{\alpha}({\rm i}u)^{\alpha}.$ (3.5)
Assumption 2 implies that the limit of $\frac{\alpha n^{\alpha}(1-a_{n})}{K\Gamma(1-\alpha)}$ exists and we denote it by $\kappa$. Consequently, for any fixed $u$, we have
$\hat{\rho}_{n}(u)\to\frac{\kappa}{\kappa+({\rm i}u)^{\alpha}},\quad\text{as}\ n\to\infty,$
and the limit happens to be the Fourier transform of $f^{\alpha,\kappa}(x)=\kappa x^{\alpha-1}E_{\alpha,\alpha}(-\kappa x^{\alpha})$. Therefore,
$F_{n}(x)=\sum_{i=1}^{\lfloor nx\rfloor}\frac{(1-a_{n})\psi^{(n)}_{i}}{a_{n}}\to\int_{0}^{x}f^{\alpha,\kappa}(s){\rm d}s,\quad\text{as}\ n\to\infty.$ (3.6)
By Dini’s theorem, this convergence is uniform.
In fact, for the density function itself, the convergence $\hat{\rho}_{n}\to\hat{f}^{\alpha,\kappa}$ implies only that the density function $\rho_{n}$ converges weakly to $f^{\alpha,\kappa}$. We now proceed to prove that this convergence also holds in $L^{2}$. To this end, we first provide some estimates for $\hat{\phi}$. Note that in $\hat{\rho}_{n}$, the function $\hat{\phi}$ appears as the form $\hat{\phi}(u/n)=:\hat{\phi}_{n}(u)$. We regard this as the Fourier transform of the step function $\phi_{n}(x)=\phi_{\lfloor nx\rfloor}$, which is consistent with the representation of $\rho_{n}$.
Lemma 2.
For any $|u|>1$, we have $|\hat{\phi}_{n}(u)|\leq|u|^{-\alpha}$. Additionally, there exist constants $c_{1},c_{2}>0$ such that
$|1-\hat{\phi}_{n}(u)|\geq\left\{\begin{aligned} c_{1}|u/n|^{\alpha},\quad\text{if}\ |u|\leq 1,\\ c_{2},\quad\text{if}\ |u|>1.\end{aligned}\right$.
Proof.
Since $\lim_{x\to\infty}\phi_{n}(x)=0$, integration by parts yields
$\begin{split}\hat{\phi}_{n}(u)=\int_{0}^{\infty}{\rm e}^{-{\rm i}ux}\phi_{n}(x){\rm d}x=\frac{1}{{\rm i}u}\sum_{k=1}^{\infty}{\rm e}^{-{\rm i}uk}\phi_{nk}.\end{split}$
Since $\|\phi\|=1$, for any $|u|>1$, we have
$|\hat{\phi}_{n}(u)|\leq\frac{1}{|u|}\sum_{k=1}^{\infty}\phi_{nk}\leq\frac{1}{|u|}\leq\frac{1}{|u|^{\alpha}}.$
From (3.5), there exists $\delta<1$ such that for all $|u|<\delta$, we have
$|1-\hat{\phi}_{n}(u)|\geq\frac{K\Gamma(1-\alpha)}{2\alpha}\left|\frac{u}{n}\right|^{\alpha}.$
Moreover, $\frac{|1-\hat{\phi}_{n}(u)|}{|u/n|^{\alpha}}$ is a continuous function on $\delta,1]$, and thus attains its minimum at some point $u_{0}\in[\delta,1]$. Suppose $|1-\hat{\phi}_{n}(u_{0})|=0$, that is $\text{Re}(\hat{\phi}_{n}(u_{0}))=\mathbb{E}[\cos u_{0}U_{1}/n]=1$. This implies that $U_{1}\in\{2kn\pi/u_{0},k=0,1,2,\ldots\}$ almost surely, but this is clearly impossible. Therefore, there exists $m=\frac{|1-\hat{\phi}(u_{0})|}{|u_{0}/n|^{\alpha}}>0$ such that
$|1-\hat{\phi}_{n}(u)|\geq m\left|\frac{u}{n}\right|^{\alpha}.$
Consequently, for $|u|\leq 1$, taking $c_{1}=\min\left\{\frac{K\Gamma(1-\alpha)}{2\alpha},m\right\}$ yields the desired result.
Similarly, since $|\hat{\phi}_{n}(u)|\leq|u|^{-\alpha}$, there exists $\delta^{\prime}>1$ such that for all $|u|>\delta^{\prime}$, we have $|1-\hat{\phi}_{n}(u)|\geq 1/2$. Moreover, $|1-\hat{\phi}_{n}(u)|$ is continuous on $1,\delta^{\prime}]$, and thus attains its minimum at some point $u_{1}\in[1,\delta^{\prime}]$. Suppose $|1-\hat{\phi}_{n}(u_{1})|=0$, that is $\text{Re}(\hat{\phi}_{n}(u_{1}))=\mathbb{E}[\cos u_{1}U_{1}/n]=1$. This implies that $U_{1}\in\{2kn\pi/u_{1},k=0,1,2,\ldots\}$ almost surely, but this is clearly impossible. Therefore, there exists $m^{\prime}=|1-\hat{\phi}_{n}(u_{1})|>0$ such that
$|1-\hat{\phi}_{n}(u)|\geq m^{\prime}.$
Consequently, for $|u|>1$, taking $c_{2}=\min\left\{1/2,m^{\prime}\right\}$ yields the desired result. ∎
Proposition 3.
The sequence $\rho_{n}$ converges to $f^{\alpha,\kappa}$ in the sense of $L^{2}$, i.e.,
$\|\rho_{n}-f^{\alpha,\kappa}\|_{2}:=\int_{0}^{1}\left[\rho_{n}(x)-f^{\alpha,\kappa}(x)\right]^{2}{\rm d}x\to 0,\ \text{as}\ n\to\infty.$ (3.7)
Proof.
To establish the $L^{2}$-convergence, we employ the Fourier isometry
$\|\rho_{n}-f^{\alpha,\kappa}\|_{2}=\frac{1}{2\pi}\|\hat{\rho}_{n}-\hat{f}^{\alpha,\kappa}\|_{2}.$
which reduces the problem to verifying $L^{2}$-convergence of the corresponding characteristic functions. Given that $\widehat{\rho}_{n}\to\widehat{f}^{\alpha,\kappa}$ pointwise, it suffices to verify that the sequence $\widehat{\rho}_{n}$ satisfies the dominated convergence theorem.
According to estimates in Lemma 2, when $|u|\leq 1$,
$|\hat{\rho}_{n}(u)|=\left|\frac{(1-a_{n})\hat{\phi}_{n}(u)}{1-a_{n}\hat{\phi}_{n}(u)}\right|\leq\frac{|1-a_{n}|}{|1-\hat{\phi}_{n}(u)|}\leq\frac{c_{1}}{|u|^{\alpha}}.$
when $|u|>1$,
$|\hat{\rho}_{n}(u)|=\left|\frac{(1-a_{n})\hat{\phi}_{n}(u)}{1-a_{n}\hat{\phi}_{n}(u)}\right|\leq\frac{|\hat{\phi}_{n}(u)|}{|1-\hat{\phi}(u/n)|}\leq\frac{1}{c_{2}|u|^{\alpha}}.$
Thus, there exists $C=\max\{c_{1},1/c_{2}\}$ such that for all $u\in\mathbb{R}$, we have
$|\hat{\rho}_{n}(u)|\leq 1\wedge\frac{C}{|u|^{\alpha}},$
and the right-hand side
$\int_{\mathbb{R}}\left(1\wedge\frac{C}{|u|^{\alpha}}\right)^{2}{\rm d}u=2\left(\int_{0}^{C^{-\alpha}}1{\rm d}u+\int_{C^{-\alpha}}^{\infty}\frac{C^{2}}{u^{2\alpha}}{\rm d}u\right)=2C^{-\alpha}+\frac{2C^{2\alpha-1}}{1-2\alpha}<\infty$
is $L^{2}$-integrable. The result thus follows by applying the dominated convergence theorem. ∎
3.2 Proof of main theorem
We proceed with the proof following the aforementioned four steps. First, we need to demonstrate how the convergence target $\Lambda_{t}$ can be expressed in terms of the operator $\mathcal{G}^{\alpha-1}$ in (3.2).
Proposition 4.
The equality $\Lambda_{t}=(\mathcal{G}^{\alpha-1}Y)_{t}$ holds almost surely for all $t\in[0,1]$.
Proof.
Since the paths of $Y_{t}$ have $1/2-\epsilon$ Hölder continuity, for $\alpha\in(1/2,1)$, the exponent $\alpha-1$ fall within (1/2,0)(-1/2,0), which is well-defined in this context of GFO. By definition of (3.2),
$(\mathcal{G}^{\alpha-1}Y)(t)=\frac{{\rm d}}{{\rm d}t}\int_{0}^{t}(Y(s)-Y(0))f^{\alpha,\kappa}(t-s){\rm d}s.$
For $\varepsilon>0$, define the operator
$(\mathcal{G}^{\alpha}_{\varepsilon}Y)(t)=\int_{0}^{t-\varepsilon}(Y(s)-Y(0))f^{\alpha,\kappa}(t-s){\rm d}s.$
Then, for any $t\in[0,1]$, we almost surely have
[…]
This convergence is uniform since the $L^{2}$-integrability of $f^{\alpha,\kappa}$. Consequently, we obtain
$(\mathcal{G}^{\alpha-1}Y)(t)=\frac{{\rm d}}{{\rm d}t}\lim_{\varepsilon\to 0}(\mathcal{G}^{\alpha}_{\varepsilon}Y)(t)=\lim_{\varepsilon\to 0}\frac{{\rm d}}{{\rm d}t}(\mathcal{G}^{\alpha}_{\varepsilon}Y)(t)=\Lambda_{t}.$
Proposition 5.
The equality $\widetilde{\Lambda}_{t}^{(n)}=(\mathcal{G}^{\alpha-1}Y^{(n)})_{t}$ holds almost surely for all $t\in[0,1]$.
Proof.
It should be noted that $Y^{(n)}$ is actually a rescaled version of the martingale $M^{(n)}$. To ensure its paths belong to a Hölder space, we adopt the linear interpolation
$Y_{t}^{(n)}=\theta t+\frac{\theta}{n\omega_{n}}\left[(1-nt+\lfloor nt\rfloor)M_{\lfloor nt\rfloor-1}^{(n)}+(nt-\lfloor nt\rfloor)M_{\lfloor nt\rfloor}^{(n)}\right],$
which is piecewise differentiable, and for $t\in(t_{i},t_{i+1})$, it holds that
$\frac{{\rm d}Y^{(n)}(t)}{{\rm d}t}=\theta+\frac{\theta}{\omega_{n}}\Delta M_{i}^{(n)}.$
By the Assumption 3, its paths is Lipschitz continuous, the GFO is well-defined on $Y_{t}^{(n)}$. Thus, by the definition (3.2),
$\begin{split}&\int_{0}^{t}(Y^{(n)}(s)-Y^{(n)}(0))f^{\alpha,\kappa}(t-s){\rm d}s=\int_{0}^{t}\int_{0}^{t-s}f^{\alpha,\kappa}(u){\rm d}u\frac{{\rm d}(Y^{(n)}(s)-Y^{(n)}(0))}{{\rm d}s}{\rm d}s\\ &=\sum_{i=0}^{\lfloor nt\rfloor-1}\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{i}^{(n)}\right)\int_{t_{i}}^{t_{i+1}}\int_{0}^{t-s}f^{\alpha,\kappa}(u){\rm d}u{\rm d}s+\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{\lfloor nt\rfloor}^{(n)}\right)\int_{t_{\lfloor nt\rfloor}}^{t}\int_{0}^{t-s}f^{\alpha,\kappa}(u){\rm d}u{\rm d}s.\end{split}$
Thus, by the definition (3.2),
$\begin{split}&(\mathcal{G}^{\alpha-1}Y^{(n)})(t)=\frac{{\rm d}}{{\rm d}t}\int_{0}^{t}(Y^{(n)}(s)-Y^{(n)}(0))f^{\alpha,\kappa}(t-s){\rm d}s\\ &=\sum_{i=0}^{\lfloor nt\rfloor-1}\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{i}^{(n)}\right)\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s+\left(\theta+\frac{\theta}{\omega_{n}}\Delta M_{\lfloor nt\rfloor}^{(n)}\right)\int_{t_{\lfloor nt\rfloor}}^{t}f^{\alpha,\kappa}(t-s){\rm d}s.\end{split}$
The second step, we prove that $Y^{(n)}\Rightarrow Y$ in Hölder space $\mathcal{C}^{\lambda}$ for $\lambda<1/2$. The Donsker invariance principle for convergence in Hölder spaces was first established by Lamperti, (1962). Račkauskas and Suquet, (2004) provided necessary and sufficient conditions for the convergence of i.i.d. partial sum processes in Hölder spaces, while results for martingale difference sequences were derived by Giraudo, (2016) from a dynamical systems perspective.
Proposition 6.
For any $\lambda<1/2$, $Y^{(n)}$ converges weakly to $Y$ in Hölder space $\mathcal{C}^{\lambda}([0,1])$.
Proof.
Note that
$Y_{t}^{(n)}=\theta t+\frac{\theta}{n\omega_{n}}\left(\sum_{i=0}^{\lfloor nt\rfloor-1}\Delta M^{(n)}_{i}+(nt-\lfloor nt\rfloor)\Delta M^{(n)}_{\lfloor nt\rfloor}\right)$,
by Theorem 2.2 in Giraudo, (2016), it suffices to prove that
$\lim_{t\to\infty}t^{p}\mathbb{P}\left(|\Delta M_{k}^{(n)}|>t\right)=0\ \text{and}\ \mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2}|\mathcal{F}_{k-1}\right]\in L^{p/2}.$ (3.8)
Then
$n^{-1/2}\left(\sum_{i=0}^{\lfloor nt\rfloor-1}\Delta M^{(n)}_{i}+(nt-\lfloor nt\rfloor)\Delta M^{(n)}_{\lfloor nt\rfloor}\right)\Rightarrow\sigma W_{t}\ \text{in}\ \mathcal{C}^{1/2-1/p}([0,1]),$
where $\sigma^{2}=\lim_{n\to\infty}\frac{1}{n}\mathbb{E}[\sum_{i=0}^{n-1}\Delta M^{(n)}_{i}]^{2}\in L^{1}$. For $p^{\prime}>0$, if $\mathbb{E}|\Delta M_{k}^{(n)}|^{p^{\prime}}<\infty$, by Markov’s inequality, for any positive $p<p^{\prime}$, we have
$t^{p}\mathbb{P}\left(|\Delta M_{k}^{(n)}|>t\right)\leq t^{p-p^{\prime}}\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p^{\prime}}\to 0,\text{as}\ t\to\infty.$
By Jensen’s inequality,
$\mathbb{E}\left(\mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2}|\mathcal{F}_{k-1}\right]\right)^{\frac{p}{2}}\leq\mathbb{E}\left(\mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2\cdot\frac{p}{2}}|\mathcal{F}_{k-1}\right]\right)=\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p}<\infty.$
Therefore, Assumption 3 ensures that condition (3.8) holds for any $p>2$. Moreover,
$\lim_{n\to\infty}\frac{1}{n}\mathbb{E}\left[\sum_{i=0}^{n-1}\Delta M^{(n)}_{i}\right]^{2}=\lim_{n\to\infty}\sum_{i=0}^{n-1}\frac{1}{n}\mathbb{E}\left[\mathbb{E}\left(|\Delta M_{i}^{(n)}|^{2}|\mathcal{F}_{i-1}\right)\right]=\lim_{n\to\infty}\sum_{i=0}^{n-1}\frac{1}{n}\mathbb{E}\left[U^{2}(\lambda_{i}^{(n)})\right]=\gamma^{2},$
together with $\frac{\theta}{\sqrt{n}\omega}\to\nu$, we have the weak convergence
$Y_{t}^{(n)}\Rightarrow\theta t+\nu\gamma W_{t}=Y_{t}$
in $\mathcal{C}^{1/2-1/p}([0,1])$, for any $p>2$. ∎
Subsequently, we establish the continuity of the operator $\mathcal{G}^{\alpha-1}$, thereby enabling the application of the continuous mapping theorem. Indeed, this result is encompassed within Proposition 2.2 of Horvath et al., (2024), where a proof of continuity for general GFOs is provided. We directly apply their result to obtain the following proposition.
Proposition 7.
For any $\alpha\in(1/2,1)$, there exists $\lambda\in(1-\alpha,1/2)$ such that the operator $\mathcal{G}^{\alpha}$ is continuous from $\mathcal{C}^{\lambda}([0,1])$ to $\mathcal{C}^{\lambda+\alpha-1}([0,1])$.
This continuity together with the fact $Y^{(n)}\Rightarrow Y$, we can obtain that $\mathcal{G}^{\alpha-1}Y^{(n)}\Rightarrow\mathcal{G}^{\alpha-1}Y$ by continuous mapping theorem. From the above representations, this entails
$\widetilde{\Lambda}_{t}^{(n)}\Rightarrow\Lambda_{t}\ \text{in}\ \mathcal{C}^{\lambda+\alpha-1}([0,1]).$
According to Lemma 3.10 in Horvath et al., (2024), we can readily extend this weak convergence to the continuous function space $\mathcal{C}([0,1])$ and Skorokhod space $\mathcal{D}([0,1])$.
We have established the convergence of $\widetilde{\Lambda}_{t}^{(n)}$. However, our primary process of interest is $\Lambda_{t}^{(n)}$ given by (2.5), with $\widetilde{\Lambda}_{t}^{(n)}$ serving merely as an auxiliary process. Finally, we will demonstrate that the two processes are asymptotically indistinguishable. By virtue of Theorem 3.1 in Billingsley, (2013), it follows that $\Lambda_{t}^{(n)}\Rightarrow\Lambda_{t}$.
Proposition 8.
The two processes $\Lambda_{t}^{(n)}$ and $\widetilde{\Lambda}_{t}^{(n)}$ are asymptotically indistinguishable in $\mathcal{C}([0,1])$, that is,
$\sup_{t\in[0,1]}\left|\Lambda_{t}^{(n)}-\widetilde{\Lambda}_{t}^{(n)}\right|\Rightarrow 0.$
Proof.
By comparing (2.5) with (3.4), we have
$\begin{split}&\left|\Lambda_{t}^{(n)}-\widetilde{\Lambda}_{t}^{(n)}\right|\\ &\leq\left|(1-a_{n})\theta\right|+\theta\left|\sum_{i=1}^{\lfloor nt\rfloor}(1-a_{n})\psi_{i}^{(n)}-\int_{0}^{t}f^{\alpha,\kappa}(u){\rm d}u\right|+\frac{\theta}{\omega_{n}}\left|\Delta M_{\lfloor nt\rfloor}^{(n)}\int_{t_{\lfloor nt\rfloor}}^{t}f^{\alpha,\kappa}(t-s){\rm d}s\right|\\ &+\frac{\theta}{n\omega_{n}}\left|\sum_{i=0}^{\lfloor nt\rfloor-1}\left[n(1-a_{n})\psi_{\lfloor nt\rfloor-i}^{(n)}-n\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s\right]\Delta M_{i}^{(n)}\right|=I_{1}+I_{2}+I_{3}+I_{4}.\end{split}$
It is evident that $I_{1}\to 0$. By (3.6) and Dini’s theorem, $I_{2}$ converges to zero uniformly. As for $I_{3}$,
$\mathbb{E}|\sup_{t\in[0,1]}I_{3}|\leq\frac{\theta}{n\omega_{n}}\mathbb{E}\left|\Delta M_{1}^{(n)}\right|\sup_{s\in(0,t-t_{\lfloor nt\rfloor}]}\left|f^{\alpha,\kappa}(s)\right|\leq\frac{C}{n\omega_{n}}\left(\frac{1}{n}\right)^{\alpha-1}\sim n^{1/2-\alpha}\to 0.$
Finally, we focus on $I_{4}$, aiming to prove that its supremum converges to zero in probability. Noting that $n(1-a_{n})\psi_{\lfloor nt\rfloor}^{(n)}/\rho_{n}(t)=a_{n}\to 1$, we will henceforth replace $n(1-a_{n})\psi_{\lfloor nt\rfloor-i}^{(n)}$ with $\rho_{n}(t-t_{i})$. Then, we employ the left-endpoint rule to approximate the integral $\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s$, and this approximation is uniformly since $f^{\alpha,\kappa}\in L^{r}$ for $r<1/(1-\alpha)$.
Therefore, for any $\varepsilon>0$,
$\begin{split}\mathbb{P}\left(\sup_{t\in[0,1]}|I_{4}|\geq\varepsilon\right)&\sim\mathbb{P}\left(\sup_{k=1,\ldots,n}\left|\frac{\theta}{n\omega_{n}}\sum_{i=0}^{k-1}\left[\rho_{n}\left(t_{k}-t_{i}\right)-f^{\alpha,\kappa}(t_{k}-t_{i})\right]\Delta M_{i}^{(n)}\right|\geq\varepsilon\right)\\ &\leq\left(\frac{\theta}{\varepsilon n\omega_{n}}\right)^{2}\mathbb{E}\left|\sum_{i=0}^{n-1}\left[\rho_{n}(t_{k}-t_{i})-f^{\alpha,\kappa}(t_{k}-t_{i})\right]\Delta M_{i}^{(n)}\right|^{2}\\ &=\left(\frac{\theta}{\varepsilon n\omega_{n}}\right)^{2}\sum_{i=0}^{n-1}\left[\rho_{n}(t_{k}-t_{i})-f^{\alpha,\kappa}(t_{k}-t_{i})\right]^{2}\mathbb{E}\left[\Delta M_{i}^{(n)}\right]^{2}\\ &=\frac{\theta^{2}\sup_{n}\mathbb{E}\left[\Delta M_{i}^{(n)}\right]^{2}}{\varepsilon^{2}n\omega_{n}^{2}}\sum_{i=1}^{n}\left[\rho_{n}(t_{i})-f^{\alpha,\kappa}(t_{i})\right]^{2}\frac{1}{n}\\ &\lesssim C_{\varepsilon}\|\rho_{n}-f^{\alpha,\kappa}\|_{2}^{2}\end{split}$
By Proposition 3, it follows that the above probability converges to zero. Consequently, $\Lambda_{t}^{(n)}$ and $\widetilde{\Lambda}_{t}^{(n)}$ are asymptotically indistinguishable. ∎
4 Specific examples for volatility models
4.1 Gamma-GED-EGARCH($\infty$) model
When we use the score-driven model to characterize volatility, we focus on the following observed time series:
$y_{n}=\sqrt{\varphi(\lambda_{n})}\varepsilon_{n},\quad\varepsilon_{n}|\mathcal{F}_{n-1}\stackrel{{\scriptstyle d}}{{\sim}}\text{density}\ f(\cdot).$
If $\varepsilon_{n}$ has unit variance, $\sqrt{\varphi(\lambda_{n})}$ is in fact the conditional volatility of $y_{n}$. Here, the function $\varphi:\mathbb{R}\to\mathbb{R}^{+}$ is monotonic and differentiable, referred to as the link function. Moreover, if $\varphi$ is the identity mapping (restricted to $\mathbb{R}^{+}$), then the time-varying parameter $\lambda_{n}$ directly characterizes the conditional variance of $y_{n}$. According to Buccheri et al., (2021), the score in this case can be expressed as
$\nabla_{n}=\frac{-\varphi^{\prime}(\lambda_{n})}{2\varphi(\lambda_{n})}\left[1+\frac{f^{\prime}\left(\frac{y_{n}}{\sqrt{\varphi(\lambda_{n})}}\right)}{f\left(\frac{y_{n}}{\sqrt{\varphi(\lambda_{n})}}\right)}\frac{y_{n}}{\sqrt{\varphi(\lambda_{n})}}\right].$
In this section, we set $\varphi(x)={\rm e}^{2x}$, which implies that $\lambda_{n}:=\ln\sigma_{n}$ is the log-volatility. We assume that $f$ is the density of Generalized Error Distribution (GED) with parameter $\nu$, that is
$f(x)=\frac{1}{2^{1+1/\nu}\Gamma(1+1/\nu)}\exp\left(-\frac{|x|^{\nu}}{2}\right).$
In this case, the score is given by
$\nabla_{n}=\frac{\nu}{2}\left|\frac{y_{n}}{{\rm e}^{\lambda_{n}}}\right|^{\nu}-1=\frac{\nu}{2}|\varepsilon_{n}|^{\nu}-1.$
Since the Fisher information is a constant, we set $S(\lambda_{n})=1$. Thus, the dynamic of $\lambda_{n}$ is given by
$\lambda_{n}=\omega+\sum_{i=1}^{n}\phi_{i}\left[\lambda_{n-i}+\frac{\nu}{2}|\varepsilon_{n-i}|^{\nu}-1\right].$
Following the terminology in Harvey and Lange, (2017), we refer to this model as the Gamma-GED-EGARCH($\infty$). When $\nu=2$, the GED reduces to standard normal distribution, and this model corresponds to the EGARCH model in Nelson, (1991) but without asymmetry. Based on the previous definitions, we define the following sequence scaling process:
$\begin{split}&X_{t}^{(n)}=\frac{1}{\sqrt{n}}\sum_{i=1}^{\lfloor nt\rfloor}\left[\exp(\lambda_{i}^{(n)})\right]^{\frac{1-a_{n}}{\omega_{n}}\theta}\varepsilon_{i},\quad\varepsilon_{i}|\mathcal{F}_{i-1}\stackrel{{\scriptstyle d}}{{\sim}}\ \text{GED}(\nu),\\ &\Lambda_{t}^{(n)}:=\frac{1-a_{n}}{\omega_{n}}\theta\lambda^{(n)}_{\lfloor nt\rfloor},\quad\lambda_{t}^{(n)}=\omega_{n}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\frac{\nu}{2}|\varepsilon_{n-i}|^{\nu}-1\right],\end{split}$ (4.1)
where $\phi_{i}\sim K/i^{1+\alpha}$, $1-a_{n}\sim n^{-\alpha}$, $\omega_{n}\sim n^{-1/2}$.
Theorem 9.
For any $\alpha\in(1/2,1)$, as $n\to\infty$, the pair $(X_{t}^{(n)},\Lambda_{t}^{(n)})$ defined in (4.1) converges weakly to $(X_{t},\Lambda_{t})$ in the Skorohod space,
$\begin{split}&X_{t}=\int_{0}^{t}{\rm e}^{\Lambda_{s}}{\rm d}B_{s},\\ &\Lambda_{t}=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\sqrt{\nu}\kappa\mu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}.\end{split}$ (4.2)
Where $B_{t},W_{t}$ are independent Brownian motions, $\kappa=\lim_{n\to\infty}\frac{\alpha n^{\alpha}(1-a_{n})}{K\Gamma(1-\alpha)}$, $\mu=\lim_{n\to\infty}\frac{\theta}{\sqrt{n}\omega_{n}}$.
Proof.
Denote $\mathbb{E}_{k-1}[\cdot]:=\mathbb{E}[\cdot|\mathcal{F}_{k-1}]$. By the property of GED, for any $p\geq 2$,
$\mathbb{E}_{k-1}[|\varepsilon_{k}|^{p}]=\frac{2^{p/\nu}}{\nu\Gamma(1+1/\nu)}\Gamma\left(\frac{p+1}{\nu}\right)<\infty.$
Thus,
$\mathbb{E}[|\nabla_{k}|^{p}]=\sum_{i=0}^{p}\binom{k}{i}(-1)^{k-i}\left(\nu/2\right)^{i}\mathbb{E}[|\varepsilon_{k}|^{\nu i}]<\infty.$
Especially, it can be easily computed that
$U^{2}(\lambda_{k})=\mathbb{E}_{k-1}[\nabla_{k}^{2}]=\frac{\nu^{2}}{4}\mathbb{E}_{k-1}[|\varepsilon|^{2\nu}]-\nu\mathbb{E}_{k-1}[|\varepsilon|^{\nu}]+1=\nu.$
The second equation then follows from Theorem 1. From the fact that $\varepsilon$ with its score are naturally uncorrelated, by Donsker’ invariance principle,
$\left(\sum_{i=1}^{\lfloor nt\rfloor}\frac{\varepsilon_{i}}{\sqrt{n}},\sum_{i=1}^{\lfloor nt\rfloor}\frac{\nabla_{i}}{\sqrt{n}}\right)\Rightarrow(B_{t},W_{t})$
are two independent Brownian motions. ∎
4.2 A extension as rough volatility approximations
The limit (4.2) is actually a rough volatility model, where $\Lambda_{t}$ characterizes the log-volatility of assets log-price $X_{t}$. Note that $\Lambda_{t}$ follows a rough OU process but is different with that in Gatheral et al., (2018) which characterizes log-volatility with a OU process driven by rough Brownian motion. Because in this volatility process (4.2), the kernel $(t-s)^{\alpha-1}$ not only emerge in noise but also in drift term. This setting is analogous with rough CIR in Jaisson and Rosenbaum, (2016) or rough Heston in El Euch and Rosenbaum, (2019).
However, before it can be considered a “useful” rough volatility model, there are still two important issues we need to deal with. First, the initial value of $(X_{t},\Lambda_{t})$ is zero in (4.2); second, the two Brownian motions are independent, which prevents the model from capturing the leverage effect in the market.
For the first one, we adopt the strategy in Wang and Cui, (2025), add a baseline in the dynamic of $\lambda_{t}^{(n)}$, that is
$\lambda_{t}^{(n)}=\omega_{t}^{(n)}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\frac{\nu}{2}|\varepsilon_{t-i}|^{\nu}-1\right],$
where
$\omega_{t}^{(n)}=\omega_{n}+\xi\omega_{n}\left(\frac{1}{1-a_{n}}\left(1-\sum_{s=1}^{t-1}\phi_{s}^{(n)}\right)-\sum_{s=1}^{t-1}\phi_{s}^{(n)}\right),$
with $\xi>0$. It result the limit $\Lambda_{t}$ becomes
$\Lambda_{t}=\theta\xi+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\sqrt{\nu}\kappa\mu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}.$
[…]
$g(x)=\frac{(\rho+\zeta)^{2}}{2}|x|{\rm e}^{-\rho x-\zeta|x|}.$
In the framework of qausi-score driven model, it recovers the asymmetric structure analogous with the EGARCH model, that is
$\lambda_{t}^{(n)}=\omega_{n}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\rho\varepsilon_{t-i}+\zeta\left(|\varepsilon_{t-i}|-\sqrt{2/\pi}\right)\right].$
In summary, we consider the following sequence scaling qausi-score driven long memory volatility model,
$\begin{split}&X_{t}^{(n)}=X_{0}-\frac{1}{2n}\sum_{i=1}^{\lfloor nt\rfloor}\left[\exp(2\lambda_{i}^{(n)})\right]^{\frac{1-a_{n}}{\omega_{n}}\theta}+\frac{1}{\sqrt{n}}\sum_{i=1}^{\lfloor nt\rfloor}\left[\exp(\lambda_{i}^{(n)})\right]^{\frac{1-a_{n}}{\omega_{n}}\theta}\varepsilon_{i},\quad\varepsilon_{i}|\mathcal{F}_{i-1}\stackrel{{\scriptstyle d}}{{\sim}}\ \mathcal{N}(0,1),\\ &\Lambda_{t}^{(n)}:=\frac{1-a_{n}}{\omega_{n}}\theta\lambda^{(n)}_{\lfloor nt\rfloor},\quad\lambda_{t}^{(n)}=\omega_{t}^{(n)}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\rho\varepsilon_{t-i}+\zeta\left(|\varepsilon_{t-i}|-\sqrt{2/\pi}\right)\right].\end{split}$ (4.3)
It can be easily verified that the limit of (4.3) is the following rough volatility model:
$\begin{split}&X_{t}=X_{0}-\frac{1}{2}\int_{0}^{t}{\rm e}^{2\Lambda_{s}}{\rm d}s+\int_{0}^{t}{\rm e}^{\Lambda_{s}}{\rm d}B_{s},\\ &\Lambda_{t}=\theta\xi+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\sqrt{\rho^{2}+\zeta^{2}}\kappa\mu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s},\\ &\text{Cov}(B_{t},W_{t})=\rho t.\end{split}$ (4.4)
4.3 Numerical simulation
Based on the convergence result, we will simulate the time series (4.3) to approximate rough volatility model (4.4). Accordingly, the simulated paths can be employed to price a variety of options via Monte Carlo methods. In this part, the parameter choices are aligned with those adopted in prior numerical studies, such as Callegaro et al., (2021), Ma and Wu, (2022) and Wang and Cui, (2025), these values are:
$\begin{split}&X_{0}=\log 100,\quad\Lambda_{0}=\log\sqrt{0.0392},\quad\rho=-0.681,\\ &\kappa=0.1,\quad\theta=\log\sqrt{0.3156},\quad\sqrt{\rho^{2}+\zeta^{2}}\mu=0.331,\quad\alpha=0.62.\end{split}$
Note that the value of $\Lambda_{0}$ and $\theta$ implies that volatility is begin with $\sqrt{0.0392}\approx 19.80\%$, and the long trem level is $\sqrt{0.3156}\approx 56.18\%$. To visualize the volatility behavior, we first simulate Equation (4.3) over a time horizon $T=5$ with $n=1000$, varying the roughness parameter $\alpha$. Figure 1 shows that the path of the volatility $\sigma_{t}=\exp(\Lambda_{t})$ by simulating (4.3) for different $\alpha$.
Figure 1: Volatility paths for different $\alpha$ values
It can be observed that as $\alpha$ increases, the volatility paths exhibit greater smoothness, and the mean-reverting behavior becomes more evident. In other words, a smaller $\alpha$ not only implies rougher paths but also corresponds to a slower mean-reversion, which is consistent with the observations reported in Gatheral et al., (2018).
We proceed to price a range of path-independent and path-dependent options, including European, Asian, Lookback, and Barrier options, using Monte Carlo simulations. The Algorithm 1 outlines the computational steps for simulating asset paths and pricing these options.
Algorithm 1 Monte Carlo Option Pricing with EGARCH($\infty$) Approximation
Input: Model parameters ($\alpha$, $\kappa$, $\theta$, $\mu$, $\rho$), initial value ($X_{0}$, $\Lambda_{0}$), number of simulated paths $M$, time steps per year $n$, time horizon $T$, strike price $K$, Up-In barrier $B_{u}$, Down-Out barrier $B_{d}$.
Output: Option prices and standard errors.
Initialization: Compute discrete model parameters:
$\phi_{1}=1-\frac{1}{\Gamma(1-\alpha)\cdot 2^{\alpha}},\quad\phi_{i}=\frac{1}{\Gamma(1-\alpha)}\left(\frac{1}{i^{\alpha}}-\frac{1}{(i+1)^{\alpha}}\right)\text{ for }i\geq 2.$
$\omega_{n}=\frac{\theta}{\mu\sqrt{n}}$, $a_{n}=1-\kappa n^{-\alpha}$, $\xi=\frac{\Lambda_{0}}{\theta}$.
Key step: Generate $M$ asset price paths $S_{t}={\rm e}^{X_{t}}$ by simulating Equation (4.3).
for each path do
  Compute various option payoffs $V_{i}$:
  • European: $\max(S_{T}-K,0)$ for call, $\max(K-S_{T},0)$ for put
  • Asian: $\max(\frac{1}{nT}\sum_{i=1}^{nT}S_{t_{i}}-K,0)$ for call, $\max(K-\frac{1}{nT}\sum_{i=1}^{nT}S_{t_{i}},0)$ for put
  • Lookback: $\max(\max_{i=0,\ldots,nT}S_{t_{i}}-K,0)$ for call, $\max(K-\min_{i=0,\ldots,nT}S_{t_{i}},0)$ for put
  • Barrier: $\max(S_{T}-K,0)1_{\max S_{t_{i}}\geq B_{u}}$ for call, $\max(K-S_{T},0)1_{\min S_{t_{i}}>B_{d}}$ for put
end for
return option price estimates $V=\frac{1}{M}\sum_{i=1}^{M}V_{i}$, and standard errors $\sqrt{\frac{\sum_{i=1}^{M}(V_{i}-V)^{2}}{(M-1)M}}$.
Table 1 reports the option prices computed via Algorithm 1, with standard errors shown in parentheses. We set $M=10^{5}$, $n=500$, $T=1$, with strike prices $K=80,90,100,110,120$. For the barrier options, the barrier level is set to $B_{u}=110$ for the Up-In call and $B_{d}=90$ for the Down-Out put.
Table 1: Option pricing results under the rough volatility model (4.4)
Strike European Asian Lookback Barrier
Call Put Call Put Call Put Call Put
80 21.4331 1.4731 20.0771 0.1216 37.2152 2.6954 18.5647 0.0000
(0.0613) (0.0135) (0.0371) (0.0027) (0.0477) (0.0173) (0.0667) (0.0000)
90 13.9924 4.0324 11.1154 1.1600 27.2152 7.3847 13.0543 0.0000
(0.0537) (0.0237) (0.0328) (0.0097) (0.0477) (0.0280) (0.0552) (0.0000)
100 8.4271 8.4671 4.7219 4.7665 17.2152 15.4721 8.2995 0.1694
(0.0439) (0.0350) (0.0237) (0.0205) (0.0477) (0.0341) (0.0440) (0.0030)
110 4.7040 14.7440 1.5043 11.5489 9.4057 25.4721 4.7040 1.0715
(0.0336) (0.0454) (0.0137) (0.0299) (0.0419) (0.0341) (0.0336) (0.0103)
120 2.4492 22.4892 0.3638 20.4084 4.8067 35.4721 2.4492 2.8810
(0.0244) (0.0537) (0.0066) (0.0352) (0.0321) (0.0341) (0.0244) (0.0204)
To accelerate the path generation process, we employ two techniques: parallel computing and FFT-based convolution. Parallel computing distributes the simulation of independent paths across multiple processors, achieving near-linear speedup. The FFT-based approach reformulates the convolution operation in Equation (4.3) as:
$\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\rho\varepsilon_{t-i}+\zeta\left(|\varepsilon_{t-i}|-\sqrt{2/\pi}\right)\right]=\left(\Phi^{(n)}*\Psi^{(n)}\right)_{t},$
where $\Phi^{(n)}:=(\phi_{i}^{(n)})_{i=1,\ldots,nT}$ and $\Psi^{(n)}:=\left(\lambda_{i}^{(n)}+\rho\varepsilon_{i}+\zeta\left(|\varepsilon_{i}|-\sqrt{2/\pi}\right)\right)_{i=0,\ldots,nT-1}$. By applying Fast Fourier Transform (FFT), the computational complexity is reduced from $\mathcal{O}((nT)^{2})$ to $\mathcal{O}(nT\log nT)$, significantly improving efficiency for large $n$ and $T$. Table 2 compares the average computational time for generating a single path with and without FFT acceleration, demonstrating substantial performance gains as $n$ increases.
Table 2: Computation time with and without FFT acceleration (in seconds)
n(T=1)n\ (T=1) 100 250 500 1000 2500 5000
With FFT 0.0029 0.0073 0.0163 0.0383 0.2952 1.0789
Without FFT 0.0104 0.0586 0.2333 0.9208 5.8149 23.1063
5 Conclusion
This study reveals the connection between long memory score-driven models and the rough OU process, the former offering statistical robustness and the latter exhibiting rich dynamic behavior, especially in volatility. Specifically, we find that when $\alpha\in(1/2,1)$, the coefficients of the lag terms decay at a power-law rate of order $1+\alpha$, and the limiting process is driven by a fractional Brownian motion with Hurst parameter $H=\alpha-1/2$, which implies roughness. This result is inspired by the work of Jaisson and Rosenbaum, (2016) on scaling limits of Hawkes processes. In our proof, however, we introduce a novel approach based on GFOs. Leveraging their continuity in Hölder spaces, we directly establish the convergence of the parameter process in the corresponding Hölder space, leading to a more direct argument.
We apply our theoretical results to volatility modeling. When the time-varying parameter governed by the score-driven mechanism represents the log-volatility, the model converges to a new rough volatility model in which the log-volatility follows a rough OU process. More precisely, it is a process obtained by applying a Riemann-Liouville fractional operator of order $\alpha$ to an OU process, which differs from that in Gatheral et al., (2018).
In the numerical experiments, we approximate this rough volatility model using an EGARCH($\infty$) specification. As a quasi score-driven model, it incorporates the score of asymmetric distributions can thus capture the leverage effect. In this sense, our results may be seen as a long memory extension of Nelson, (1990)’s result for AR(1)(1) Exponential ARCH. The proposed Monte Carlo algorithm, optimized through FFT-based convolution and parallel computing, achieves significant computational efficiency, making it suitable for pricing diverse options, including European, Asian, Lookback, and Barrier options.
The results of this paper can be further developed in future work. In fact, the derivation of the OU-type limiting process relies on the assumption that the second-order conditional moment of the score, denoted by $U$, is asymptotically constant. While this condition is not difficult to satisfy—for instance, it always holds when the scaling function $S$ is chosen as the 1/2-1/2 power of the Fisher information—we believe that the results can be extended to more general forms of $U$, yielding corresponding limits of the form (2.8). Such an extension would require new techniques, as the GFO approach used here would break down: the limit process in this case cannot be represented as a GFO applied to the associated diffusion.
Acknowledgements
The authors express gratitude to Yingli Wang for insightful discussions and valuable suggestions, which significantly improved the quality of this work.

### v2

@@
 Building on this framework, we carry out numerical simulations and option pricing analyses, offering new tools for rough volatility modeling and simulation.
-Keywords: Score-driven models, Rough volatility, Weak convergence, Option pricing
-2020 MSC: 60F17, 60G22, 62M10, 91B70
+keywords:
+Score-driven models , Rough volatility , Weak convergence , Option pricing
+2020 MSC:
+60F17 , 60G22 , 62M10 , 91B70
+journal: Journal of Econometrics
+[…]
+[label1]organization=School of Mathematics, Shanghai University of Finance and Economics,city=Shanghai, postcode=200433, country=China
 1 Introduction
-Score-driven models, also known as generalized autoregressive score models, were proposed by Creal et al., (2013) and Harvey, (2013) as a class of observation-driven time series models in which parameter updates are driven by the score—the gradient of the log-likelihood function.
-By directly linking the score of the data distribution to parameter dynamics in a unified framework, these models inherit flexibility and generality, encompassing many classical models such as GARCH (Bollerslev, , 1986) and ACD (Engle and Russell, , 1998) models.
-Theoretically, score-driven models have been proven optimal in minimizing the Kullback-Leibler divergence between the true distribution and the postulated distribution (Gorgi et al., , 2024).
+Score-driven models, also known as generalized autoregressive score models, were proposed by creal2013generalized and harvey2013dynamic as a class of observation-driven time series models in which parameter updates are driven by the score-the gradient of the log-likelihood function.
+By directly linking the score of the data distribution to parameter dynamics in a unified framework, these models inherit flexibility and generality, encompassing many classical models such as GARCH (bollerslev1986generalized) and ACD (engle1998autoregressive) models.
+Theoretically, score-driven models have been proven optimal in minimizing the Kullback-Leibler divergence between the true distribution and the postulated distribution (gorgi2023optimality).
 Moreover, in statistical inference, the model’s closed-form likelihood function facilitates efficient maximum likelihood estimation, exhibiting lower computational costs compared to parameter-driven models.
 In application, they also demonstrate robust empirical performance across diverse fields such as economics, finance and biology.
-Within just a few years, the literature on SD models has grown to nearly 400 and can be found online at http://www.gasmodel.com/gaspapers.htm.
-As a remarkably influential model in time series analysis, it naturally raises the question: what would its continuous-time counterpart look like?
-Buccheri et al., (2021) have investigated its continuous-time limit, obtaining a two-dimensional diffusion process.
-Similarly, Wu and He, (2024) investigated the continuous-time limit of the quasi score-driven volatility model proposed by Blasques et al., (2023), and derived a correlated stochastic volatility model.
-However, the score-driven models they studied only incorporate first-order or finite-order lags—these are all Markov processes.
-In fact, when the model was first introduced in Creal et al., (2013), the question was raised as to whether it could be extended to an infinite-lag form.
-Janus et al., (2014) and Opschoor and Lucas, (2019) pursued such extensions by proposing long memory score-driven models.
+Within just a few years, the literature on SD models has grown to nearly 400 and can be found online at gasmodel.com.
+As a remarkably influential model in time series analysis, one naturally raises the question: what would its continuous-time counterpart look like?
+Building on the method of nelson1990arch for deriving the continuous-time limit of GARCH, buccheri2021continuous have investigated continuous-time limit of score-driven volatility models, obtaining a two-dimensional diffusion process.
+Similarly, wu2024continuous investigated the continuous-time limit of the quasi score-driven volatility models proposed by blasques2023quasi, and derived a correlated stochastic volatility model.
+However, the score-driven models they studied only incorporate first-order or finite-order lags, which are all Markovian.
+In fact, when the model was first introduced in creal2013generalized, the question was raised as to whether it could be extended to an infinite-lag form.
+janus2014long and opschoor2019fractional pursued such extensions by proposing long memory score-driven models.
 This naturally leads to the question: What would be the continuous-time counterpart of such a long memory version?
-On the other hand, since Gatheral et al., (2018) discovered that volatility is rough, rough volatility models have gained popularity in financial modeling.
-These models are characterized by the use of fractional Brownian motion with a Hurst parameter $H<1/2$, whose paths are rougher than those of Brownian motion—hence the name.
-El Euch et al., (2018) investigated the microstructure of rough volatility by employing Hawkes processes to model the arrival dynamics in limit order books.
-This theoretical basis originates from the discovery of Jaisson and Rosenbaum, (2016) that the scaling limit of nearly unstable heavy-tailed Hawkes processes is a rough Cox-Ingersoll-Ross (CIR) process.
+On the other hand, since gatheral2018volatility discovered that volatility is rough, rough volatility models have gained popularity in financial modeling.
+These models are characterized by the use of fractional Brownian motion with a Hurst parameter $H<1/2$, whose paths are rougher than those of Brownian motion, hence the name.
+el2018microstructural investigated the microstructure of rough volatility by employing Hawkes processes to model the arrival dynamics in limit order books.
+This theoretical basis originates from the discovery of jaisson2016rough that the scaling limit of nearly unstable heavy-tailed Hawkes processes is a rough Cox-Ingersoll-Ross (CIR) process.
 It is worth noting that Hawkes processes are self-exciting (similar to observation-driven mechanisms) and exhibit long memory, which provides inspiration for us.
-In particular, Cai et al., (2024) studied the discrete-time counterpart of the Hawkes process—the INAR($\infty$) process—and obtained similar convergence results.
+In particular, cai2024scaling studied the INAR($\infty$) process, a discrete-time counterpart of the Hawkes process, and obtained similar convergence results.
 This motivates our investigation into the scaling limit of score-driven models when extended to incorporate long memory.
@@
 Consequently, long memory score-driven volatility models can serve as discrete-time approximations for such rough volatility models, facilitating efficient Monte Carlo simulations for financial applications.
-To the best of our knowledge, apart from Wang and Cui, (2025)’s approach using INAR($\infty$) processes to approximate the rough Heston model, this study presents one of the few methodologies employing time series models to approximate rough volatility processes.
+To the best of our knowledge, apart from wang2025rough’s approach using INAR($\infty$) processes to approximate the rough Heston model, this study presents one of the few methodologies employing time series models to approximate rough volatility processes.
 Moreover, we present specific examples of long memory score-driven volatility models to further demonstrate the practical versatility and broad applicability of our theoretical results.
@@
 Section 2 introduces the long memory score-driven models that we focus on and provides a heuristic derivation of their continuous-time limits.
-In Section 3, we establish the main convergence results and presents their proofs.
-Section 4 applies these results to volatility modeling, illustrating with the Gamma-GED-EGARCH($\infty$) model and its extension.
+In Section 3, we present our main convergence results.
+Proofs of the theorems are given in the main text when brief; all remaining proofs are deferred to the A.
+Section 4 applies these results to volatility modeling, illustrating with the Gamma-GED-EGARCH($\infty$) model and its extension, with the related proofs provided in B.
 We conducts corresponding approximations of rough volatility models and option pricing.
@@
 2 Long memory score-driven models
-Since Engle, (1982) discovered that economic data exhibit conditional heteroskedasticity, it has gradually been recognized that certain characteristics of time series are not immutable.
-Specifically, let $\{y_{n}\}_{n\in\mathbb{N}}$ be a time series and $\mathcal{F}{n}=\sigma(y_{n},y_{n-1},\ldots,y_{0})$ be the $\sigma$-algebra it generates.
+Since engle1982autoregressive discovered that economic data exhibit conditional heteroskedasticity, it has gradually been recognized that certain characteristics of time series are not immutable.
+Specifically, let $\{y_{n}\}_{n\in\mathbb{N}}$ be a time series and $\mathcal{F}_{n}=\sigma(y_{n},y_{n-1},\ldots,y_{0})$ be the $\sigma$-algebra it generates.
 The conditional density
@@
 with a time-varying parameter $\lambda_{n}$.
-To address this issue, Cox et al., (1981) proposed two modeling approaches: observation-driven models, where the dynamics of the time-varying parameters are governed by past observations of the series itself, and parameter-driven models, in which the parameter dynamics are driven by another stochastic process.
+To address this issue, cox1981statistical proposed two modeling approaches: observation-driven models, where the dynamics of the time-varying parameters are governed by past observations of the series itself, and parameter-driven models, in which the parameter dynamics are driven by another stochastic process.
 The score-driven model is a class of observation-driven models whose key idea is to update time-varying parameters based on the score of the observation.
 Specifically, the parameter update follows the form:
-$\lambda_{n+1}=\omega+\beta\lambda_{n}+\alpha S(\lambda_{n})\nabla_{n}.$ (2.1)
+$\lambda_{n+1}=\omega+\beta\lambda_{n}+\alpha S(\lambda_{n})\nabla_{n},$ (2.1)
 where $\nabla_{n}=\frac{\partial\log p(y_{n}|\lambda_{n},\mathcal{F}{n-1})}{\partial\lambda_{n}}$ denotes the partial derivative of the log-density function of observations with respect to the parameter, commonly referred to as the (Fisher) score in statistics.
@@
 The function $S(\cdot)$ is known as the scaling function, which accounts for the curvature of the log-density function.
-It is typically chosen as the negative exponent of the conditional Fisher information, i.e., $S(\lambda_{n})=[\mathbb{E}(\nabla_{n}^{2}|\mathcal{F}{n-1})]^{-a}$, with common choices for $a$ being 0, $1/2$, or $1$.
+It is typically chosen as the negative exponent of the conditional Fisher information, i.e., $S(\lambda_{n})=[\mathbb{E}(\nabla_{n}^{2}|\mathcal{F}_{n-1})]^{-a}$, with common choices for $a$ being 0, $1/2$, or $1$.
 To illustrate, when the density function $p(\cdot)$ is chosen to be that of a normal distribution and the parameter $\lambda$ is identified as the conditional variance $\sigma^{2}$, with $a=1$, the score-driven model reduces to the classical GARCH model.
 The equation given by (2.1) represents a first-order score-driven model, as the relationship between the score and the parameter is lagged by only one period.
-This framework can be extended to higher orders, such as $p$-th order or even infinite order, as in Janus et al., (2014),
+This framework can be extended to higher orders, such as $p$-th order or even infinite order, as in janus2014long,
 $\lambda_{n}=\omega+\sum_{i=1}^{n}\phi_{i}S(\lambda_{n-i})\nabla_{n-i}.$
-Noting that our time index starts from 0, the summation at time $n$ includes only $n$ terms.
-The term “infinite” refers to the dependence on the entire history.
+It should be noted that our time index starts from zero; thus, the term “infinite” here refers to dependence on the entire history-from time 0 up to the present.
 Additionally, we omit the autoregressive term, as it can be absorbed into the coefficients $\phi_{i}$ through the moving average representation.
@@
 $\sigma^{2}_{n}=\omega+\sum_{i=1}^{n}\phi_{i}y^{2}_{n-i}=\omega+\sum_{i=1}^{n}\phi_{i}\sigma^{2}_{n-i}+\sum_{i=1}^{n}\phi_{i}(y^{2}_{n-i}-\sigma^{2}_{n-i}),$
-which includes an additional infinite autoregressive component. To address this, Opschoor and Lucas, (2019) proposed the following formulation:
+which includes an additional infinite autoregressive component. To address this, opschoor2019fractional proposed the following formulation:
 $\lambda_{n}=\omega+\sum_{i=1}^{n}\phi_{i}[\lambda_{n-i}+S(\lambda_{n-i})\nabla_{n-i}].$ (2.2)
@@
 This case is referred to as nearly unstable.
-Specifically, we set $\phi_{i}\sim K/i^{1+\alpha}$, exhibiting power-law decay, where $\alpha\in(1/2,1)$, a form referred to as “heavy-tailed” in Jaisson and Rosenbaum, (2016).
+Specifically, we set $\phi_{i}\sim K/i^{1+\alpha}$, exhibiting power-law decay, where $\alpha\in(1/2,1)$, a form referred to as “heavy-tailed” in jaisson2016rough.
 This condition is critical for the limit to generate roughness.
@@
 where $\{W_{t}\}_{t\geq 0}$ is a standard Brownian motion.
-And according to Cai et al., (2024), in that case $1/2<\alpha<1$, when $1-a_{n}\sim n^{-\alpha}$ tends to zero and the coefficients $\phi_{i}\sim K/i^{1+\alpha}$ exhibit power-law decay, the following weak convergence holds:
+And according to cai2024scaling, in that case $1/2<\alpha<1$, when $1-a_{n}\sim n^{-\alpha}$ tends to zero and the coefficients $\phi_{i}\sim K/i^{1+\alpha}$ exhibit power-law decay, the following weak convergence holds:
 $n(1-a_{n})\psi^{(n)}_{\lfloor nx\rfloor}\to f^{\alpha,\kappa}(x):=\kappa x^{\alpha-1}E_{\alpha,\alpha}(-\kappa x^{\alpha}),\quad\mbox{as}\ n\to\infty,$ (2.6)
@@
 $\Lambda_{t}=\theta\int_{0}^{t}f^{\alpha,\kappa}(s){\rm d}s+\nu\int_{0}^{t}f^{\alpha,\kappa}(t-s)U(\Lambda_{s}){\rm d}W_{s}.$ (2.7)
-In fact, similar to Proposition 4.10 in El Euch et al., (2018), the SDE (2.7) has the same solution as the SDE
+In fact, similar to Proposition 4.10 in el2018microstructural, the SDE (2.7) has the same solution as the SDE
 $\Lambda_{t}=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\kappa\nu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}U(\Lambda_{s}){\rm d}W_{s}.$ (2.8)
 Since $\alpha-1\in(-1/2,0)$, the term $\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}$ corresponds to a fractional Brownian motion with Hurst parameter $H<1/2$, which means (2.8) falls into the rough case.
-It is worth mentioning that when $U(x)=\sqrt{x}$, this actually corresponds to the rough fractional CIR process described in Jaisson and Rosenbaum, (2016).
+It is worth mentioning that when $U(x)=\sqrt{x}$, this actually corresponds to the rough fractional CIR process described in jaisson2016rough.
 3 Main results
@@
 Assumption 1.
-There exists a constant $\gamma<\infty$, such that $U(x)\to\gamma$, as $x\to\infty$.
+The function $U:\mathbb{R}\to\mathbb{R}$ is locally bounded and converges to a finite limit $\gamma$ as $|x|\to\infty$.
 Assumption 2.
@@
 Assumption 3.
-For each $n$, the martingale difference sequence $\{\Delta M_{k}^{(n)}\}_{k\in\mathbb{N}}$ is strictly stationary. Furthermore, for any $p\geq 2$,
-$\sup_{n\in\mathbb{N}}\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p}<\infty.$
+The martingale difference arrays $\{\Delta M_{k}^{(n)}\}_{k\leq k_{n},n\geq 1}$ satisfy: for every $p>2$,
+$\sup_{n\geq 1}\max_{k\leq k_{n}}\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p}<\infty.$
+Remark 1.
+Without loss of generality, we often set the terminal time $T=1$, so that $k_{n}=\lfloor nT\rfloor=n$ is a common and sufficient choice.
+This assumption controls all higher-order moments to ensure the validity of the following theorem for all $\alpha\in(1/2,1)$.
+In fact, if only a finite $p$-th moment exists for some $p>2$, the theorem still holds for $\alpha\in[\frac{p+2}{2p},1)$.
 Then, we have the following theorem.
@@
 In fact, we restrict our attention to the case of rough OU process in (2.8).
-This is because our proof strategy fundamentally differs from that of Jaisson and Rosenbaum, (2016) or Wang and Cui, (2025): while they show that the Hawkes process (equivalent to the integrated intensity in probability) converges to a integrated rough CIR process, we directly establish the convergence of the parameter process itself to a rough process, rather than its integral.
+This is because our proof strategy fundamentally differs from that of jaisson2016rough or wang2025rough: while they show that the Hawkes process (equivalent to the integrated intensity in probability) converges to a integrated rough CIR process, we directly establish the convergence of the parameter process itself to a rough process, rather than its integral.
 The difficulty in dealing with the process itself lies in verifying its tightness and we attempt to circumvent this by applying the continuous mapping theorem.
-To this end, we regard the rough diffusion process as the image of a diffusion process under the generalized fractional operator (GFO), which is inspired by Horvath et al., (2024).
+To this end, we regard the rough diffusion process as the image of a diffusion process under the generalized fractional operator (GFO), which is inspired by horvath2024functional.
 Specifically, for $\lambda\in(0,1)$ and $\alpha\in(-\lambda,1-\lambda)$, $\mathcal{G}^{\alpha}$ is the GFO associated to kernel
@@
 $(\mathcal{G}^{\alpha}f)(t)=\left\{\begin{aligned} &\int_{0}^{t}(f(s)-f(0))\frac{{\rm d}}{{\rm d}t}g(t-s){\rm d}s,\quad\text{if}\ \alpha\geq 0,\\ &\frac{{\rm d}}{{\rm d}t}\int_{0}^{t}(f(s)-f(0))g(t-s){\rm d}s,\quad\text{if}\ \alpha<0.\end{aligned}\right$.
-When $g(x)=x^{\alpha}$, this reduces to the classical Riemann-Liouville operator.
+(3.2)
+When $g(x)=\frac{x^{\alpha}}{\Gamma(\alpha+1)}$, this reduces to the classical Riemann-Liouville operator.
 In this paper, we choose $g(x)=f^{\alpha,\kappa}(x)\in\mathcal{L}^{\alpha-1}$ and focus on the case $\alpha\in(1/2,1)$.
@@
 Define $Y_{t}:=\theta t+\nu\gamma W_{t}$, and
-$\Lambda_{t}=\theta\int_{0}^{t}f^{\alpha,\kappa}(s){\rm d}s+\nu\gamma\int_{0}^{t}f^{\alpha,\kappa}(t-s){\rm d}W_{s}$
-is actually has the same solution as the (3.1). Consider the partition $\mathcal{T}:=\{t_{i}=i/n,i=0,1,\ldots,n\}$. We then define
+$\Lambda_{t}:=\theta\int_{0}^{t}f^{\alpha,\kappa}(s){\rm d}s+\nu\gamma\int_{0}^{t}f^{\alpha,\kappa}(t-s){\rm d}W_{s}$ (3.4)
+actually is the solution of the equation (3.1) we aim to converge to:
+Proposition 2.
+The process $\Lambda$ is the unique strong solution to the following rough SDE:
+$\Lambda_{t}=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\kappa(\theta-\Lambda_{s}){\rm d}s+\frac{\gamma\kappa\nu}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}{\rm d}W_{s}.$
+Proof.
+See A.1. ∎
+Accordingly, in the subsequent proof, we take the clearer form of equation (3.4) as the target of convergence. Consider the partition $\mathcal{T}:=\{t_{i}=i/n,i=0,1,\ldots,n\}$. We then define
 $Y_{t}^{(n)}=\theta t+\frac{\theta}{n\omega_{n}}\left[(1-nt+\lfloor nt\rfloor)M_{\lfloor nt\rfloor-1}^{(n)}+(nt-\lfloor nt\rfloor)M_{\lfloor nt\rfloor}^{(n)}\right],$ (3.5)
@@
 3.1 Converges to $f^{\alpha,\kappa}(x)$
-Jaisson and Rosenbaum, (2016) prove that the mean of geometric sums converges in distribution to a random variable with density function $f^{\alpha,\kappa}(x)$.
+jaisson2016rough prove that the mean of geometric sums converges in distribution to a random variable with density function $f^{\alpha,\kappa}(x)$.
 Our setting differs slightly because the summed terms follow a discrete distribution.
@@
 where $\hat{\phi}$ denotes the Fourier transform of $\phi$, or equivalently, the characteristic function of the random variable $U_{1}$.
-Since $\phi_{i}\sim\frac{K}{i^{1+\alpha}}$, by the Karamata-Tauberian theorem (see Cai et al., (2024)), we can characterize the behavior of $\hat{\phi}$ near the origin,
+Since $\phi_{i}\sim\frac{K}{i^{1+\alpha}}$, by the Karamata-Tauberian theorem (see cai2024scaling), we can characterize the behavior of $\hat{\phi}$ near the origin,
 $1-\hat{\phi}(u)\underset{u\to 0}{\sim}\frac{K\Gamma(1-\alpha)}{\alpha}({\rm i}u)^{\alpha}.$ (3.7)
@@
 Proof.
-Since $\lim_{x\to\infty}\phi_{n}(x)=0$, integration by parts yields
-$\begin{split}\hat{\phi}_{n}(u)=\int_{0}^{\infty}{\rm e}^{-{\rm i}ux}\phi_{n}(x){\rm d}x=\frac{1}{{\rm i}u}\sum_{k=1}^{\infty}{\rm e}^{-{\rm i}uk}\phi_{nk}.\end{split}$
-Since $\|\phi\|=1$, for any $|u|>1$, we have
-$|\hat{\phi}_{n}(u)|\leq\frac{1}{|u|}\sum_{k=1}^{\infty}\phi_{nk}\leq\frac{1}{|u|}\leq\frac{1}{|u|^{\alpha}}.$
-From (3.5), there exists $\delta<1$ such that for all $|u|<\delta$, we have
-$|1-\hat{\phi}_{n}(u)|\geq\frac{K\Gamma(1-\alpha)}{2\alpha}\left|\frac{u}{n}\right|^{\alpha}.$
-Moreover, $\frac{|1-\hat{\phi}_{n}(u)|}{|u/n|^{\alpha}}$ is a continuous function on $\delta,1]$, and thus attains its minimum at some point $u_{0}\in[\delta,1]$.
-Suppose $|1-\hat{\phi}_{n}(u_{0})|=0$, that is $\text{Re}(\hat{\phi}_{n}(u_{0}))=\mathbb{E}[\cos u_{0}U_{1}/n]=1$.
-This implies that $U_{1}\in\{2kn\pi/u_{0},k=0,1,2,\ldots\}$ almost surely, but this is clearly impossible.
-Therefore, there exists $m=\frac{|1-\hat{\phi}(u_{0})|}{|u_{0}/n|^{\alpha}}>0$ such that
-$|1-\hat{\phi}_{n}(u)|\geq m\left|\frac{u}{n}\right|^{\alpha}.$
-Consequently, for $|u|\leq 1$, taking $c_{1}=\min\left\{\frac{K\Gamma(1-\alpha)}{2\alpha},m\right\}$ yields the desired result.
-Similarly, since $|\hat{\phi}_{n}(u)|\leq|u|^{-\alpha}$, there exists $\delta^{\prime}>1$ such that for all $|u|>\delta^{\prime}$, we have $|1-\hat{\phi}_{n}(u)|\geq 1/2$.
-Moreover, $|1-\hat{\phi}_{n}(u)|$ is continuous on $1,\delta^{\prime}]$, and thus attains its minimum at some point $u_{1}\in[1,\delta^{\prime}]$.
-Suppose $|1-\hat{\phi}_{n}(u_{1})|=0$, that is $\text{Re}(\hat{\phi}_{n}(u_{1}))=\mathbb{E}[\cos u_{1}U_{1}/n]=1$.
-This implies that $U_{1}\in\{2kn\pi/u_{1},k=0,1,2,\ldots\}$ almost surely, but this is clearly impossible.
-Therefore, there exists $m^{\prime}=|1-\hat{\phi}_{n}(u_{1})|>0$ such that
-$|1-\hat{\phi}_{n}(u)|\geq m^{\prime}.$
-Consequently, for $|u|>1$, taking $c_{2}=\min\left\{1/2,m^{\prime}\right\}$ yields the desired result. ∎
+See A.2. ∎
+Utilizing the Fourier isometry, the estimates for the characteristic function $\hat{\phi}_{n}$ established in Lemma 3 help us analyze the $L^{2}$ distance $\|\rho_{n}-f^{\alpha,\kappa}\|_{2}$ in the Fourier domain.
+We are now equipped to obtain the stronger convergence result.
 Proposition 4.
 The sequence $\rho_{n}$ converges to $f^{\alpha,\kappa}$ in the sense of $L^{2}$, i.e.,
-$\|\rho_{n}-f^{\alpha,\kappa}\|_{2}:=\int_{0}^{1}\left[\rho_{n}(x)-f^{\alpha,\kappa}(x)\right]^{2}{\rm d}x\to 0,\ \text{as}\ n\to\infty.$ (3.7)
+$\|\rho_{n}-f^{\alpha,\kappa}\|_{2}:=\left(\int_{0}^{1}\left[\rho_{n}(x)-f^{\alpha,\kappa}(x)\right]^{2}{\rm d}x\right)^{1/2}\to 0,\ \text{as}\ n\to\infty.$ (3.9)
 Proof.
-To establish the $L^{2}$-convergence, we employ the Fourier isometry
-$\|\rho_{n}-f^{\alpha,\kappa}\|_{2}=\frac{1}{2\pi}\|\hat{\rho}_{n}-\hat{f}^{\alpha,\kappa}\|_{2}.$
-which reduces the problem to verifying $L^{2}$-convergence of the corresponding characteristic functions.
-Given that $\widehat{\rho}_{n}\to\widehat{f}^{\alpha,\kappa}$ pointwise, it suffices to verify that the sequence $\widehat{\rho}_{n}$ satisfies the dominated convergence theorem.
-According to estimates in Lemma 2, when $|u|\leq 1$,
-$|\hat{\rho}_{n}(u)|=\left|\frac{(1-a_{n})\hat{\phi}_{n}(u)}{1-a_{n}\hat{\phi}_{n}(u)}\right|\leq\frac{|1-a_{n}|}{|1-\hat{\phi}_{n}(u)|}\leq\frac{c_{1}}{|u|^{\alpha}}.$
-when $|u|>1$,
-$|\hat{\rho}_{n}(u)|=\left|\frac{(1-a_{n})\hat{\phi}_{n}(u)}{1-a_{n}\hat{\phi}_{n}(u)}\right|\leq\frac{|\hat{\phi}_{n}(u)|}{|1-\hat{\phi}(u/n)|}\leq\frac{1}{c_{2}|u|^{\alpha}}.$
-Thus, there exists $C=\max\{c_{1},1/c_{2}\}$ such that for all $u\in\mathbb{R}$, we have
-$|\hat{\rho}_{n}(u)|\leq 1\wedge\frac{C}{|u|^{\alpha}},$
-and the right-hand side
-$\int_{\mathbb{R}}\left(1\wedge\frac{C}{|u|^{\alpha}}\right)^{2}{\rm d}u=2\left(\int_{0}^{C^{-\alpha}}1{\rm d}u+\int_{C^{-\alpha}}^{\infty}\frac{C^{2}}{u^{2\alpha}}{\rm d}u\right)=2C^{-\alpha}+\frac{2C^{2\alpha-1}}{1-2\alpha}<\infty$
-is $L^{2}$-integrable. The result thus follows by applying the dominated convergence theorem. ∎
+See A.3. ∎
 3.2 Proof of main theorem
 We proceed with the proof following the aforementioned four steps.
-First, we need to demonstrate how the convergence target $\Lambda_{t}$ can be expressed in terms of the operator $\mathcal{G}^{\alpha-1}$ in (3.2).
+First, we need to demonstrate how the convergence target $\Lambda_{t}$ can be expressed in terms of the operator $\mathcal{G}^{\alpha-1}$ defined in (3.3).
 Proposition 5.
@@
 The second step, we prove that $Y^{(n)}\Rightarrow Y$ in Hölder space $\mathcal{C}^{\lambda}$ for $\lambda<1/2$.
-The Donsker invariance principle for convergence in Hölder spaces was first established by Lamperti, (1962).
-Račkauskas and Suquet, (2004) provided necessary and sufficient conditions for the convergence of i.i.d.
-partial sum processes in Hölder spaces, while results for martingale difference sequences were derived by Giraudo, (2016) from a dynamical systems perspective.
+The Donsker invariance principle for convergence in Hölder spaces was first established by lamperti1962convergence.
+ravckauskas2004necessary provided necessary and sufficient conditions for the convergence of i.i.d.
+partial sum processes in Hölder spaces, while results for martingale difference sequences were derived by giraudo2016holderian from a dynamical systems perspective.
+However, these results concern only the scaling limits of a single sequence and do not directly apply to triangular arrays.
+Actually, FCLT for martingale difference arrays is well established in the Skorokhod space (hall2014martingale).
+Consequently, once tightness in the Hölder space is verified, the desired convergence in this space follows immediately.
 Proposition 7.
@@
 Proof.
-Note that
-$Y_{t}^{(n)}=\theta t+\frac{\theta}{n\omega_{n}}\left(\sum_{i=0}^{\lfloor nt\rfloor-1}\Delta M^{(n)}_{i}+(nt-\lfloor nt\rfloor)\Delta M^{(n)}_{\lfloor nt\rfloor}\right)$,
-by Theorem 2.2 in Giraudo, (2016), it suffices to prove that
-$\lim_{t\to\infty}t^{p}\mathbb{P}\left(|\Delta M_{k}^{(n)}|>t\right)=0\ \text{and}\ \mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2}|\mathcal{F}_{k-1}\right]\in L^{p/2}.$ (3.8)
-Then
-$n^{-1/2}\left(\sum_{i=0}^{\lfloor nt\rfloor-1}\Delta M^{(n)}_{i}+(nt-\lfloor nt\rfloor)\Delta M^{(n)}_{\lfloor nt\rfloor}\right)\Rightarrow\sigma W_{t}\ \text{in}\ \mathcal{C}^{1/2-1/p}([0,1]),$
-where $\sigma^{2}=\lim_{n\to\infty}\frac{1}{n}\mathbb{E}[\sum_{i=0}^{n-1}\Delta M^{(n)}_{i}]^{2}\in L^{1}$.
-For $p^{\prime}>0$, if $\mathbb{E}|\Delta M_{k}^{(n)}|^{p^{\prime}}<\infty$, by Markov’s inequality, for any positive $p<p^{\prime}$, we have
-$t^{p}\mathbb{P}\left(|\Delta M_{k}^{(n)}|>t\right)\leq t^{p-p^{\prime}}\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p^{\prime}}\to 0,\text{as}\ t\to\infty.$
-By Jensen’s inequality,
-$\mathbb{E}\left(\mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2}|\mathcal{F}_{k-1}\right]\right)^{\frac{p}{2}}\leq\mathbb{E}\left(\mathbb{E}\left[|\Delta M_{k}^{(n)}|^{2\cdot\frac{p}{2}}|\mathcal{F}_{k-1}\right]\right)=\mathbb{E}\left|\Delta M_{k}^{(n)}\right|^{p}<\infty.$
-Therefore, Assumption 3 ensures that condition (3.8) holds for any $p>2$. Moreover,
-$\lim_{n\to\infty}\frac{1}{n}\mathbb{E}\left[\sum_{i=0}^{n-1}\Delta M^{(n)}_{i}\right]^{2}=\lim_{n\to\infty}\sum_{i=0}^{n-1}\frac{1}{n}\mathbb{E}\left[\mathbb{E}\left(|\Delta M_{i}^{(n)}|^{2}|\mathcal{F}_{i-1}\right)\right]=\lim_{n\to\infty}\sum_{i=0}^{n-1}\frac{1}{n}\mathbb{E}\left[U^{2}(\lambda_{i}^{(n)})\right]=\gamma^{2},$
-together with $\frac{\theta}{\sqrt{n}\omega}\to\nu$, we have the weak convergence
-$Y_{t}^{(n)}\Rightarrow\theta t+\nu\gamma W_{t}=Y_{t}$
-in $\mathcal{C}^{1/2-1/p}([0,1])$, for any $p>2$. ∎
+See A.4
 Subsequently, we establish the continuity of the operator $\mathcal{G}^{\alpha-1}$, thereby enabling the application of the continuous mapping theorem.
-Indeed, this result is encompassed within Proposition 2.2 of Horvath et al., (2024), where a proof of continuity for general GFOs is provided.
+Indeed, this result is encompassed within Proposition 2.2 of horvath2024functional, where a proof of continuity for general GFOs is provided.
 We directly apply their result to obtain the following proposition.
@@
 $\widetilde{\Lambda}_{t}^{(n)}\Rightarrow\Lambda_{t}\ \text{in}\ \mathcal{C}^{\lambda+\alpha-1}([0,1]).$
-According to Lemma 3.10 in Horvath et al., (2024), we can readily extend this weak convergence to the continuous function space $\mathcal{C}([0,1])$ and Skorokhod space $\mathcal{D}([0,1])$.
+According to Lemma 3.10 in horvath2024functional, we can readily extend this weak convergence to the continuous function space $\mathcal{C}([0,1])$ and Skorokhod space $\mathcal{D}([0,1])$.
 We have established the convergence of $\widetilde{\Lambda}_{t}^{(n)}$.
@@
 Finally, we will demonstrate that the two processes are asymptotically indistinguishable.
-By virtue of Theorem 3.1 in Billingsley, (2013), it follows that $\Lambda_{t}^{(n)}\Rightarrow\Lambda_{t}$.
+By virtue of Theorem 3.1 in billingsley2013convergence, it follows that $\Lambda_{t}^{(n)}\Rightarrow\Lambda_{t}$.
 Proposition 9.
@@
 Proof.
-By comparing (2.5) with (3.4), we have
-$\begin{split}&\left|\Lambda_{t}^{(n)}-\widetilde{\Lambda}_{t}^{(n)}\right|\\ &\leq\left|(1-a_{n})\theta\right|+\theta\left|\sum_{i=1}^{\lfloor nt\rfloor}(1-a_{n})\psi_{i}^{(n)}-\int_{0}^{t}f^{\alpha,\kappa}(u){\rm d}u\right|+\frac{\theta}{\omega_{n}}\left|\Delta M_{\lfloor nt\rfloor}^{(n)}\int_{t_{\lfloor nt\rfloor}}^{t}f^{\alpha,\kappa}(t-s){\rm d}s\right|\\ &+\frac{\theta}{n\omega_{n}}\left|\sum_{i=0}^{\lfloor nt\rfloor-1}\left[n(1-a_{n})\psi_{\lfloor nt\rfloor-i}^{(n)}-n\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s\right]\Delta M_{i}^{(n)}\right|=I_{1}+I_{2}+I_{3}+I_{4}.\end{split}$
-It is evident that $I_{1}\to 0$. By (3.6) and Dini’s theorem, $I_{2}$ converges to zero uniformly. As for $I_{3}$,
-$\mathbb{E}|\sup_{t\in[0,1]}I_{3}|\leq\frac{\theta}{n\omega_{n}}\mathbb{E}\left|\Delta M_{1}^{(n)}\right|\sup_{s\in(0,t-t_{\lfloor nt\rfloor}]}\left|f^{\alpha,\kappa}(s)\right|\leq\frac{C}{n\omega_{n}}\left(\frac{1}{n}\right)^{\alpha-1}\sim n^{1/2-\alpha}\to 0.$
-Finally, we focus on $I_{4}$, aiming to prove that its supremum converges to zero in probability.
-Noting that $n(1-a_{n})\psi_{\lfloor nt\rfloor}^{(n)}/\rho_{n}(t)=a_{n}\to 1$, we will henceforth replace $n(1-a_{n})\psi_{\lfloor nt\rfloor-i}^{(n)}$ with $\rho_{n}(t-t_{i})$.
-Then, we employ the left-endpoint rule to approximate the integral $\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s$, and this approximation is uniformly since $f^{\alpha,\kappa}\in L^{r}$ for $r<1/(1-\alpha)$.
-Therefore, for any $\varepsilon>0$,
-$\begin{split}\mathbb{P}\left(\sup_{t\in[0,1]}|I_{4}|\geq\varepsilon\right)&\sim\mathbb{P}\left(\sup_{k=1,\ldots,n}\left|\frac{\theta}{n\omega_{n}}\sum_{i=0}^{k-1}\left[\rho_{n}\left(t_{k}-t_{i}\right)-f^{\alpha,\kappa}(t_{k}-t_{i})\right]\Delta M_{i}^{(n)}\right|\geq\varepsilon\right)\\ &\leq\left(\frac{\theta}{\varepsilon n\omega_{n}}\right)^{2}\mathbb{E}\left|\sum_{i=0}^{n-1}\left[\rho_{n}(t_{k}-t_{i})-f^{\alpha,\kappa}(t_{k}-t_{i})\right]\Delta M_{i}^{(n)}\right|^{2}\\ &=\left(\frac{\theta}{\varepsilon n\omega_{n}}\right)^{2}\sum_{i=0}^{n-1}\left[\rho_{n}(t_{k}-t_{i})-f^{\alpha,\kappa}(t_{k}-t_{i})\right]^{2}\mathbb{E}\left[\Delta M_{i}^{(n)}\right]^{2}\\ &=\frac{\theta^{2}\sup_{n}\mathbb{E}\left[\Delta M_{i}^{(n)}\right]^{2}}{\varepsilon^{2}n\omega_{n}^{2}}\sum_{i=1}^{n}\left[\rho_{n}(t_{i})-f^{\alpha,\kappa}(t_{i})\right]^{2}\frac{1}{n}\\ &\lesssim C_{\varepsilon}\|\rho_{n}-f^{\alpha,\kappa}\|_{2}^{2}\end{split}$
-By Proposition 3, it follows that the above probability converges to zero. Consequently, $\Lambda_{t}^{(n)}$ and $\widetilde{\Lambda}_{t}^{(n)}$ are asymptotically indistinguishable. ∎
+See A.5
 4 Specific examples for volatility models
@@
 Moreover, if $\varphi$ is the identity mapping (restricted to $\mathbb{R}^{+}$), then the time-varying parameter $\lambda_{n}$ directly characterizes the conditional variance of $y_{n}$.
-According to Buccheri et al., (2021), the score in this case can be expressed as
+According to buccheri2021continuous, the score in this case can be expressed as
 $\nabla_{n}=\frac{-\varphi^{\prime}(\lambda_{n})}{2\varphi(\lambda_{n})}\left[1+\frac{f^{\prime}\left(\frac{y_{n}}{\sqrt{\varphi(\lambda_{n})}}\right)}{f\left(\frac{y_{n}}{\sqrt{\varphi(\lambda_{n})}}\right)}\frac{y_{n}}{\sqrt{\varphi(\lambda_{n})}}\right].$
@@
 $\lambda_{n}=\omega+\sum_{i=1}^{n}\phi_{i}\left[\lambda_{n-i}+\frac{\nu}{2}|\varepsilon_{n-i}|^{\nu}-1\right].$
-Following the terminology in Harvey and Lange, (2017), we refer to this model as the Gamma-GED-EGARCH($\infty$).
-When $\nu=2$, the GED reduces to standard normal distribution, and this model corresponds to the EGARCH model in Nelson, (1991) but without asymmetry.
+Following the terminology in harvey2017volatility, we refer to this model as the Gamma-GED-EGARCH($\infty$).
+When $\nu=2$, the GED reduces to standard normal distribution, and this model corresponds to the EGARCH model in nelson1991conditional but without asymmetry.
 Based on the previous definitions, we define the following sequence scaling process:
-$\begin{split}&X_{t}^{(n)}=\frac{1}{\sqrt{n}}\sum_{i=1}^{\lfloor nt\rfloor}\left[\exp(\lambda_{i}^{(n)})\right]^{\frac{1-a_{n}}{\omega_{n}}\theta}\varepsilon_{i},\quad\varepsilon_{i}|\mathcal{F}_{i-1}\stackrel{{\scriptstyle d}}{{\sim}}\ \text{GED}(\nu),\\ &\Lambda_{t}^{(n)}:=\frac{1-a_{n}}{\omega_{n}}\theta\lambda^{(n)}_{\lfloor nt\rfloor},\quad\lambda_{t}^{(n)}=\omega_{n}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\frac{\nu}{2}|\varepsilon_{n-i}|^{\nu}-1\right],\end{split}$ (4.1)
-where $\phi_{i}\sim K/i^{1+\alpha}$, $1-a_{n}\sim n^{-\alpha}$, $\omega_{n}\sim n^{-1/2}$.
+$\begin{split}&X_{t}^{(n)}=\frac{1}{\sqrt{\chi n}}\sum_{i=1}^{\lfloor nt\rfloor}\left[\exp(\lambda_{i}^{(n)})\right]^{\frac{1-a_{n}}{\omega_{n}}\theta}\varepsilon_{i},\quad\varepsilon_{i}|\mathcal{F}_{i-1}\stackrel{{\scriptstyle d}}{{\sim}}\ \text{GED}(\nu),\\ &\Lambda_{t}^{(n)}:=\frac{1-a_{n}}{\omega_{n}}\theta\lambda^{(n)}_{\lfloor nt\rfloor},\quad\lambda_{t}^{(n)}=\omega_{n}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\frac{\nu}{2}|\varepsilon_{n-i}|^{\nu}-1\right],\end{split}$ (4.1)
+where $\phi_{i}\sim K/i^{1+\alpha}$, $1-a_{n}\sim n^{-\alpha}$, $\omega_{n}\sim n^{-1/2}$, $\chi=\mathbb{E}(\varepsilon^{2})=2^{2/\nu}\frac{\Gamma(3/\nu)}{\Gamma(1/\nu)}$.
 Theorem 10.
@@
 Proof.
-Denote $\mathbb{E}_{k-1}[\cdot]:=\mathbb{E}[\cdot|\mathcal{F}_{k-1}]$. By the property of GED, for any $p\geq 2$,
-$\mathbb{E}_{k-1}[|\varepsilon_{k}|^{p}]=\frac{2^{p/\nu}}{\nu\Gamma(1+1/\nu)}\Gamma\left(\frac{p+1}{\nu}\right)<\infty.$
-Thus,
-$\mathbb{E}[|\nabla_{k}|^{p}]=\sum_{i=0}^{p}\binom{k}{i}(-1)^{k-i}\left(\nu/2\right)^{i}\mathbb{E}[|\varepsilon_{k}|^{\nu i}]<\infty.$
-Especially, it can be easily computed that
-$U^{2}(\lambda_{k})=\mathbb{E}_{k-1}[\nabla_{k}^{2}]=\frac{\nu^{2}}{4}\mathbb{E}_{k-1}[|\varepsilon|^{2\nu}]-\nu\mathbb{E}_{k-1}[|\varepsilon|^{\nu}]+1=\nu.$
-The second equation then follows from Theorem 1. From the fact that $\varepsilon$ with its score are naturally uncorrelated, by Donsker’ invariance principle,
-$\left(\sum_{i=1}^{\lfloor nt\rfloor}\frac{\varepsilon_{i}}{\sqrt{n}},\sum_{i=1}^{\lfloor nt\rfloor}\frac{\nabla_{i}}{\sqrt{n}}\right)\Rightarrow(B_{t},W_{t})$
-are two independent Brownian motions. ∎
+See B.1
+The limit (4.2) is actually a rough volatility model, where $\Lambda_{t}$ characterizes the log-volatility of assets log-price $X_{t}$.
+Note that $\Lambda_{t}$ follows a rough OU process but is different with that in gatheral2018volatility which characterizes log-volatility with a OU process driven by rough fractional Brownian motion.
+Because the kernel $(t-s)^{\alpha-1}$ not only emerge in noise but also in drift term, this specification is analogous with rough CIR in jaisson2016rough or rough Heston in el2019characteristic.
 4.2 A extension as rough volatility approximations
-The limit (4.2) is actually a rough volatility model, where $\Lambda_{t}$ characterizes the log-volatility of assets log-price $X_{t}$.
-Note that $\Lambda_{t}$ follows a rough OU process but is different with that in Gatheral et al., (2018) which characterizes log-volatility with a OU process driven by rough Brownian motion.
-Because in this volatility process (4.2), the kernel $(t-s)^{\alpha-1}$ not only emerge in noise but also in drift term.
-This setting is analogous with rough CIR in Jaisson and Rosenbaum, (2016) or rough Heston in El Euch and Rosenbaum, (2019).
-However, before it can be considered a “useful” rough volatility model, there are still two important issues we need to deal with.
+Before (4.2) can be considered a “useful” rough volatility model, there are still two important issues we need to deal with.
 First, the initial value of $(X_{t},\Lambda_{t})$ is zero in (4.2); second, the two Brownian motions are independent, which prevents the model from capturing the leverage effect in the market.
-For the first one, we adopt the strategy in Wang and Cui, (2025), add a baseline in the dynamic of $\lambda_{t}^{(n)}$, that is
+For the first issue, we adopt the strategy in wang2025rough, add a baseline in the dynamic of $\lambda_{t}^{(n)}$, that is
 $\lambda_{t}^{(n)}=\omega_{t}^{(n)}+\sum_{i=1}^{t}\phi_{i}^{(n)}\left[\lambda_{t-i}^{(n)}+\frac{\nu}{2}|\varepsilon_{t-i}|^{\nu}-1\right],$
@@
 Accordingly, the simulated paths can be employed to price a variety of options via Monte Carlo methods.
-In this part, the parameter choices are aligned with those adopted in prior numerical studies, such as Callegaro et al., (2021), Ma and Wu, (2022) and Wang and Cui, (2025), these values are:
+In this part, the parameter choices are aligned with those adopted in prior numerical studies, such as callegaro2021fast, ma2022fast and wang2025rough, these values are:
 $\begin{split}&X_{0}=\log 100,\quad\Lambda_{0}=\log\sqrt{0.0392},\quad\rho=-0.681,\\ &\kappa=0.1,\quad\theta=\log\sqrt{0.3156},\quad\sqrt{\rho^{2}+\zeta^{2}}\mu=0.331,\quad\alpha=0.62.\end{split}$
@@
 It can be observed that as $\alpha$ increases, the volatility paths exhibit greater smoothness, and the mean-reverting behavior becomes more evident.
-In other words, a smaller $\alpha$ not only implies rougher paths but also corresponds to a slower mean-reversion, which is consistent with the observations reported in Gatheral et al., (2018).
+In other words, a smaller $\alpha$ not only implies rougher paths but also corresponds to a slower mean-reversion, which is consistent with the observations reported in gatheral2018volatility.
 We proceed to price a range of path-independent and path-dependent options, including European, Asian, Lookback, and Barrier options, using Monte Carlo simulations.
@@
 return option price estimates $V=\frac{1}{M}\sum_{i=1}^{M}V_{i}$, and standard errors $\sqrt{\frac{\sum_{i=1}^{M}(V_{i}-V)^{2}}{(M-1)M}}$.
-Table 1 reports the option prices computed via Algorithm 1, with standard errors shown in parentheses.
-We set $M=10^{5}$, $n=500$, $T=1$, with strike prices $K=80,90,100,110,120$.
+Table 1 reports the option prices computed via Algorithm 1 for different values of the roughness parameter $\alpha$, with standard errors shown in parentheses.
+We set $M=5\times 10^{5}$, $n=500$, $T=1$, with strike prices $K=80,90,100,110,120$.
 For the barrier options, the barrier level is set to $B_{u}=110$ for the Up-In call and $B_{d}=90$ for the Down-Out put.
-Table 1: Option pricing results under the rough volatility model (4.4)
+It can be seen that as $\alpha$ decreases, European calls and puts prices rise, reflecting the fat-tailed return distributions generated by rough volatility processes.
+Asian options follow similar patterns but with attenuated sensitivity, as price averaging dampens the impact of path roughness.
+Lookback options show the most pronounced positive sensitivity to roughness, as their payoffs directly depend on extreme price realizations that become more frequent under rough volatility.
+Barrier options exhibit the most intricate response.
+While barrier calls increase with roughness, corresponding barrier puts decrease in value.
+This occurs because roughness simultaneously increases potential terminal payoffs but also raises the probability of early barrier triggering.
+Table 1: Option pricing results under the rough volatility model (4.4) for different $\alpha$
+Panel A: $\alpha=0.999$
 Strike European Asian Lookback Barrier
 Call Put Call Put Call Put Call Put
-80 21.4331 1.4731 20.0771 0.1216 37.2152 2.6954 18.5647 0.0000
+80 21.3756 1.3852 20.0912 0.1055 36.9770 2.5257 18.4050 0.0000
 (0.0270) (0.0058) (0.0163) (0.0011) (0.0210) (0.0074) (0.0295) (0.0000)
-90 13.9924 4.0324 11.1154 1.1600 27.2152 7.3847 13.0543 0.0000
+90 13.8728 3.8825 11.0856 1.0999 26.9770 7.1114 12.9005 0.0000
 (0.0236) (0.0103) (0.0145) (0.0042) (0.0210) (0.0122) (0.0243) (0.0000)
-100 8.4271 8.4671 4.7219 4.7665 17.2152 15.4721 8.2995 0.1694
+100 8.2723 8.2820 4.6574 4.6717 16.9770 15.1612 8.1390 0.1835
 (0.0192) (0.0154) (0.0104) (0.0090) (0.0210) (0.0150) (0.0193) (0.0014)
-110 4.7040 14.7440 1.5043 11.5489 9.4057 25.4721 4.7040 1.0715
+110 4.5555 14.5651 1.4513 11.4656 9.1915 25.1612 4.5555 1.1421
 (0.0146) (0.0200) (0.0059) (0.0132) (0.0183) (0.0150) (0.0146) (0.0048)
-120 2.4492 22.4892 0.3638 20.4084 4.8067 35.4721 2.4492 2.8810
+120 2.3334 22.3430 0.3423 20.3566 4.6381 35.1612 2.3334 3.0491
 (0.0105) (0.0237) (0.0028) (0.0155) (0.0140) (0.0150) (0.0105) (0.0094)
+Panel B: $\alpha=0.62$
+Strike European Asian Lookback Barrier
+Call Put Call Put Call Put Call Put
+80 21.4739 1.4614 20.1284 0.1205 37.2465 2.6685 18.6106 0.0000
+(0.0274) (0.0060) (0.0166) (0.0012) (0.0213) (0.0077) (0.0298) (0.0000)
+90 14.0136 4.0010 11.1613 1.1535 27.2465 7.3286 13.0776 0.0000
+(0.0240) (0.0106) (0.0147) (0.0043) (0.0213) (0.0125) (0.0246) (0.0000)
+100 8.4253 8.4128 4.7481 4.7402 17.2465 15.4061 8.2971 0.1731
+(0.0196) (0.0156) (0.0106) (0.0091) (0.0213) (0.0152) (0.0197) (0.0013)
+110 4.6899 14.6774 1.5125 11.5046 9.4320 25.4061 4.6899 1.0888
+(0.0150) (0.0203) (0.0061) (0.0134) (0.0187) (0.0152) (0.0150) (0.0047)
+120 2.4363 22.4237 0.3641 20.3562 4.8210 35.4061 2.4363 2.9263
+(0.0109) (0.0240) (0.0029) (0.0157) (0.0144) (0.0152) (0.0109) (0.0092)
+Panel C: $\alpha=0.501$
+Strike European Asian Lookback Barrier
+Call Put Call Put Call Put Call Put
+80 21.4982 1.4699 20.1332 0.1252 37.2660 2.6897 18.6239 0.0000
+(0.0274) (0.0060) (0.0166) (0.0012) (0.0214) (0.0077) (0.0298) (0.0000)
+90 14.0469 4.0186 11.1787 1.1707 27.2660 7.3630 13.0986 0.0000
+(0.0240) (0.0106) (0.0147) (0.0044) (0.0214) (0.0125) (0.0246) (0.0000)
+100 8.4553 8.4269 4.7727 4.7647 17.2660 15.4459 8.3241 0.1699
+(0.0196) (0.0157) (0.0106) (0.0092) (0.0214) (0.0152) (0.0197) (0.0013)
+110 4.7082 14.6799 1.5238 11.5158 9.4538 25.4459 4.7082 1.0710
+(0.0150) (0.0203) (0.0061) (0.0134) (0.0188) (0.0152) (0.0150) (0.0046)
+120 2.4448 22.4165 0.3717 20.3636 4.8332 35.4459 2.4448 2.8889
+(0.0109) (0.0240) (0.0030) (0.0158) (0.0144) (0.0152) (0.0109) (0.0091)
+To further verify the validity of our results, we compute the pricing outcomes when $\alpha=1$.
+In this case, model (4.4) reduces to the classical lognormal stochastic volatility model proposed by scott1987option:
+$\begin{split}&{\rm d}X_{t}=-\frac{1}{2}{\rm e}^{2\Lambda_{t}}{\rm d}t+{\rm e}^{\Lambda_{t}}{\rm d}B_{t},\\ &{\rm d}\Lambda_{t}=\kappa(\theta-\Lambda_{t}){\rm d}t+\sigma{\rm d}W_{t},\\ &\text{Cov}(B_{t},W_{t})=\rho t,\end{split}$
+where $\sigma=\sqrt{\rho^{2}+\zeta^{2}}\kappa\mu$.
+Using the same parameter values and Monte-Carlo settings as in the previous experiments, we report the option prices in Table 2.
+The numbers in parentheses represent the percentage errors relative to the rough volatility model with $\alpha=0.999$.
+The errors are mostly within 0.5%, with minor deviations attributable to the diminishing roughness.
+These results confirm that the option prices computed from our model are consistent with those from the classical specification.
+Table 2: Option pricing results under the lognormal stochastic volatility model (case $\alpha=1$)
+Strike European Asian Lookback Barrier
+Call Put Call Put Call Put Call Put
+80 21.3770 1.4003 20.1121 0.1086 36.9692 2.5534 18.4349 0.0000
+(-0.01%) (-1.08%) (-0.10%) (-2.85%) (0.02%) (-1.08%) (-0.16%) (0.00%)
+90 13.8839 3.9071 11.1075 1.1040 26.9692 7.1480 12.9184 0.0000
+(-0.08%) (-0.63%) (-0.20%) (-0.37%) (0.03%) (-0.51%) (-0.14%) (0.00%)
+100 8.2834 8.3067 4.6752 4.6716 16.9692 15.1966 8.1519 0.1812
+(-0.13%) (-0.30%) (-0.38%) (0.00%) (0.05%) (-0.23%) (-0.16%) (1.27%)
+110 4.5550 14.5783 1.4563 11.4527 9.1719 25.1966 4.5550 1.1266
+(0.01%) (-0.09%) (-0.34%) (0.11%) (0.21%) (-0.14%) (0.01%) (1.38%)
+120 2.3233 22.3466 0.3396 20.3360 4.6129 35.1966 2.3233 3.0172
+(0.43%) (-0.02%) (0.80%) (0.10%) (0.55%) (-0.10%) (0.43%) (1.06%)
+We then examine the shape of the implied volatility surface generated by the model, focusing specifically on the at-the-money (ATM) volatility skew $\psi(\tau)$, defined as
+$\psi(\tau):=\left|\frac{\partial\sigma_{IV}(k,\tau)}{\partial k}\right|_{k=0},$
+where $k=\log(S_{0}/K)$ denotes the log-moneyness and $\tau$ represents the time to maturity.
+Empirical evidence suggests that the volatility skew exhibits a power-law decay as time to maturity increases (gatheral2018volatility).
+Theoretical work by fukasawa2011asymptotic established that when volatility is driven by fractional Brownian motion with Hurst parameter $H$, the ATM skew asymptotically follows $\psi(\tau)\sim\tau^{H-1/2}$ for small $\tau$.
+We compute the ATM volatility skew generated from our model, as shown in Figure 2.
+To assess the power-law relationship, we employ a double logarithmic scale and perform a linear regression on the log-transformed data.
+The results largely validate a power-law behavior ($R^{2}>0.85$), with two observations consistent with prior studies: 1.
+Rougher volatility paths generate steeper initial skews that subsequently decay more rapidly as $\tau$ increases.
+This aligns with the notion that greater path roughness implies a stronger local impact of volatility shocks, resulting in a more pronounced skew effect, especially for short maturities; 2.
+In the limit case where roughness is nearly absent ($\alpha=0.999$), the volatility skew is observed to approximate a constant across maturities.
+Figure 2: Power-law relationship between ATM volatility skew and time to maturity (Log-Log plot)
+Finally, we are concerned with the computational speed of this approximation.
 To accelerate the path generation process, we employ two techniques: parallel computing and FFT-based convolution.
@@
 Specifically, we find that when $\alpha\in(1/2,1)$, the coefficients of the lag terms decay at a power-law rate of order $1+\alpha$, and the limiting process is driven by a fractional Brownian motion with Hurst parameter $H=\alpha-1/2$, which implies roughness.
-This result is inspired by the work of Jaisson and Rosenbaum, (2016) on scaling limits of Hawkes processes.
+This result is inspired by the work of jaisson2016rough on scaling limits of Hawkes processes.
 In our proof, however, we introduce a novel approach based on GFOs.
@@
 When the time-varying parameter governed by the score-driven mechanism represents the log-volatility, the model converges to a new rough volatility model in which the log-volatility follows a rough OU process.
-More precisely, it is a process obtained by applying a Riemann-Liouville fractional operator of order $\alpha$ to an OU process, which differs from that in Gatheral et al., (2018).
+More precisely, it is a process obtained by applying a Riemann-Liouville fractional operator of order $\alpha$ to an OU process, which differs from that in gatheral2018volatility.
 In the numerical experiments, we approximate this rough volatility model using an EGARCH($\infty$) specification.
 As a quasi score-driven model, it incorporates the score of asymmetric distributions can thus capture the leverage effect.
-In this sense, our results may be seen as a long memory extension of Nelson, (1990)’s result for AR(1)(1) Exponential ARCH.
+In this sense, our results may be seen as a long memory extension of nelson1990arch’s result for AR(1)(1) Exponential ARCH.
 The proposed Monte Carlo algorithm, optimized through FFT-based convolution and parallel computing, achieves significant computational efficiency, making it suitable for pricing diverse options, including European, Asian, Lookback, and Barrier options.
@@
 In fact, the derivation of the OU-type limiting process relies on the assumption that the second-order conditional moment of the score, denoted by $U$, is asymptotically constant.
-While this condition is not difficult to satisfy—for instance, it always holds when the scaling function $S$ is chosen as the 1/2-1/2 power of the Fisher information—we believe that the results can be extended to more general forms of $U$, yielding corresponding limits of the form (2.8).
+While this condition is not difficult to satisfy – for instance, it always holds when the scaling function $S$ is chosen as the 1/2-1/2 power of the Fisher information – we believe that the results can be extended to more general forms of $U$, yielding corresponding limits of the form (2.8).
 Such an extension would require new techniques, as the GFO approach used here would break down: the limit process in this case cannot be represented as a GFO applied to the associated diffusion.
@@
 The authors express gratitude to Yingli Wang for insightful discussions and valuable suggestions, which significantly improved the quality of this work.
+Appendix A The proof in Section 3
+A.1 Proof of Proposition 2
+We first introduce the notation and preliminary used throughout the proof, when the kernel is chosen as $g(x)=\frac{x^{\alpha}}{\Gamma(\alpha+1)}$ in (3.2), the operator $\mathcal{G}^{\alpha}$ becomes the Riemann-Liouville fractional integral operator of order $\alpha$.
+For $\alpha\in(1/2,1)$, we denote $I^{\alpha}=\mathcal{G}^{\alpha}$ and $D^{\alpha}=\mathcal{G}^{-\alpha}$.
+When $\alpha=1$, $I^{1}$ and $D^{1}$ represent the standard integral and differential operators, respectively.
+Furthermore, we define $F^{\alpha,\kappa}(t):=\int_{0}^{t}f^{\alpha,\kappa}(s){\rm d}s$, the cumulative distribution function of the Mittag-Leffler density $f^{\alpha,\kappa}(t)$.
+Then, the following standard but important identity holds:
+$I^{1-\alpha}f^{\alpha,\kappa}(t)=\kappa(1-F^{\alpha,\kappa}(t)),$ (A.1)
+which can be readily verified by examining the Laplace transforms of both sides.
+As Equation (3.1) is actually a linear stochastic Volterra equation, the existence and uniqueness of its strong solution are guaranteed by coutin2001stochastic.
+It therefore suffices to show that $\Lambda_{t}$ satisfies this equation.
+We apply the fractional integral operator $I^{1-\alpha}$ to $\Lambda_{t}$.
+Let $K_{t}=I^{1-\alpha}\Lambda_{t}$.
+We decompose $K_{t}$ into its deterministic and stochastic components,
+$K_{t}=K_{t}^{\text{det}}+K_{t}^{\text{stoch}}=I^{1-\alpha}\left(\theta F^{\alpha,\kappa}(t)\right)+I^{1-\alpha}\left(\nu\gamma\int_{0}^{t}f^{\alpha,\kappa}(t-s){\rm d}W_{s}\right).$
+For the deterministic part, by using identity (A.1) we obtain
+$\begin{split}K_{t}^{\text{det}}&=I^{1-\alpha}[\theta F^{\alpha,\kappa}](t)=\theta\cdot I^{1-\alpha}[I^{1}f^{\alpha,\kappa}](t)=\theta\cdot I^{1}[I^{1-\alpha}f^{\alpha,\kappa}](t)\\ &=\theta\int_{0}^{t}\left(I^{1-\alpha}f^{\alpha,\kappa}(s)\right){\rm d}s=\kappa\theta\int_{0}^{t}(1-F^{\alpha,\kappa}(s)){\rm d}s=\kappa\theta t-\kappa\theta\int_{0}^{t}F^{\alpha,\kappa}(s){\rm d}s.\end{split}$
+Similarly, combining with the stochastic Fubini theorem yields the stochastic part as follow:
+[…]
+Thus, we have
+$K_{t}=\kappa\theta t+\kappa\nu\gamma W_{t}-\kappa\int_{0}^{t}\Lambda_{s}{\rm d}s=\kappa\int_{0}^{t}(\theta-\Lambda_{s}){\rm d}s+\kappa\nu\gamma W_{t}$
+Now, we apply the fractional derivative operator $D^{1-\alpha}$ to $K_{t}$ to recover the $\Lambda_{t}$.
+According to the property $D^{1-\alpha}I^{1}=I^{\alpha}$ and the result on the action of $\mathcal{G}^{\alpha-1}$ on diffusion processes established in Proposition 5, we obtain
+[…]
+This is precisely Equation (3.1), which completes the proof.
+A.2 Proof of Lemma 3
+Since $\lim_{x\to\infty}\phi_{n}(x)=0$, integration by parts yields
+$\begin{split}\hat{\phi}_{n}(u)=\int_{0}^{\infty}{\rm e}^{-{\rm i}ux}\phi_{n}(x){\rm d}x=\frac{1}{{\rm i}u}\sum_{k=1}^{\infty}{\rm e}^{-{\rm i}uk}\phi_{nk}.\end{split}$
+Since $\|\phi\|=1$, for any $|u|>1$, we have
+$|\hat{\phi}_{n}(u)|\leq\frac{1}{|u|}\sum_{k=1}^{\infty}\phi_{nk}\leq\frac{1}{|u|}\leq\frac{1}{|u|^{\alpha}}.$
+From (3.7), there exists $\delta<1$ such that for all $|u|<\delta$, we have
+$|1-\hat{\phi}_{n}(u)|\geq\frac{K\Gamma(1-\alpha)}{2\alpha}\left|\frac{u}{n}\right|^{\alpha}.$
+Moreover, $\frac{|1-\hat{\phi}_{n}(u)|}{|u/n|^{\alpha}}$ is a continuous function on $\delta,1]$, and thus attains its minimum at some point $u_{0}\in[\delta,1]$.
+Suppose $|1-\hat{\phi}_{n}(u_{0})|=0$, that is $\text{Re}(\hat{\phi}_{n}(u_{0}))=\mathbb{E}[\cos u_{0}U_{1}/n]=1$.
+This implies that $U_{1}\in\{2kn\pi/u_{0},k=0,1,2,\ldots\}$ almost surely, but this is clearly impossible.
+Therefore, there exists $m=\frac{|1-\hat{\phi}(u_{0})|}{|u_{0}/n|^{\alpha}}>0$ such that
+$|1-\hat{\phi}_{n}(u)|\geq m\left|\frac{u}{n}\right|^{\alpha}.$
+Consequently, for $|u|\leq 1$, taking $c_{1}=\min\left\{\frac{K\Gamma(1-\alpha)}{2\alpha},m\right\}$ yields the desired result.
+Similarly, since $|\hat{\phi}_{n}(u)|\leq|u|^{-\alpha}$, there exists $\delta^{\prime}>1$ such that for all $|u|>\delta^{\prime}$, we have $|1-\hat{\phi}_{n}(u)|\geq 1/2$.
+Moreover, $|1-\hat{\phi}_{n}(u)|$ is continuous on $1,\delta^{\prime}]$, and thus attains its minimum at some point $u_{1}\in[1,\delta^{\prime}]$.
+Suppose $|1-\hat{\phi}_{n}(u_{1})|=0$, that is $\text{Re}(\hat{\phi}_{n}(u_{1}))=\mathbb{E}[\cos u_{1}U_{1}/n]=1$.
+This implies that $U_{1}\in\{2kn\pi/u_{1},k=0,1,2,\ldots\}$ almost surely, but this is clearly impossible.
+Therefore, there exists $m^{\prime}=|1-\hat{\phi}_{n}(u_{1})|>0$ such that
+$|1-\hat{\phi}_{n}(u)|\geq m^{\prime}.$
+Consequently, for $|u|>1$, taking $c_{2}=\min\left\{1/2,m^{\prime}\right\}$ yields the desired result.
+A.3 Proof of Proposition 4
+To establish the $L^{2}$-convergence, we employ the Fourier isometry
+$\|\rho_{n}-f^{\alpha,\kappa}\|_{2}=\frac{1}{2\pi}\|\hat{\rho}_{n}-\hat{f}^{\alpha,\kappa}\|_{2}.$
+which reduces the problem to verifying $L^{2}$-convergence of the corresponding characteristic functions.
+Given that $\widehat{\rho}_{n}\to\widehat{f}^{\alpha,\kappa}$ pointwise, it suffices to verify that the sequence $\widehat{\rho}_{n}$ satisfies the dominated convergence theorem.
+According to estimates in Lemma 3, when $|u|\leq 1$,
+$|\hat{\rho}_{n}(u)|=\left|\frac{(1-a_{n})\hat{\phi}_{n}(u)}{1-a_{n}\hat{\phi}_{n}(u)}\right|\leq\frac{|1-a_{n}|}{|1-\hat{\phi}_{n}(u)|}\leq\frac{c_{1}}{|u|^{\alpha}}.$
+when $|u|>1$,
+$|\hat{\rho}_{n}(u)|=\left|\frac{(1-a_{n})\hat{\phi}_{n}(u)}{1-a_{n}\hat{\phi}_{n}(u)}\right|\leq\frac{|\hat{\phi}_{n}(u)|}{|1-\hat{\phi}(u/n)|}\leq\frac{1}{c_{2}|u|^{\alpha}}.$
+Thus, there exists $C=\max\{c_{1},1/c_{2}\}$ such that for all $u\in\mathbb{R}$, we have
+$|\hat{\rho}_{n}(u)|\leq 1\wedge\frac{C}{|u|^{\alpha}},$
+and the right-hand side
+$\int_{\mathbb{R}}\left(1\wedge\frac{C}{|u|^{\alpha}}\right)^{2}{\rm d}u=2\left(\int_{0}^{C^{-\alpha}}1{\rm d}u+\int_{C^{-\alpha}}^{\infty}\frac{C^{2}}{u^{2\alpha}}{\rm d}u\right)=2C^{-\alpha}+\frac{2C^{2\alpha-1}}{1-2\alpha}<\infty$
+is $L^{2}$-integrable. The result thus follows by applying the dominated convergence theorem.
+A.4 Proof of Proposition 7
+The process under consideration is defined as
+$Y_{t}^{(n)}=\theta t+\frac{\theta}{\sqrt{n}\omega_{n}}S_{n}(t),$
+where $S_{n}(t)$ represents the normalized sum of martingale differences:
+$S_{n}(t)=\frac{1}{\sqrt{n}}\left(\sum_{i=0}^{\lfloor nt\rfloor-1}\Delta M^{(n)}_{i}+(nt-\lfloor nt\rfloor)\Delta M^{(n)}_{\lfloor nt\rfloor}\right)$.
+We establish the convergence of $S_{n}$ in the Hölder space $\mathcal{C}^{\lambda}([0,1])$ to the Brownian motion $\gamma W$ by verifying the two standard components of a functional limit theorem: the convergence of finite-dimensional distributions and tightness.
+Step 1: Convergence of finite-dimensional distributions.
+We first establish the weak convergence $S_{n}\Rightarrow\gamma W$ in the Skorokhod space $D([0,1])$, which implies the convergence of finite-dimensional distributions.
+To this end, we verify the conditions of classical FLCT for martingales (hall2014martingale).
+According to Assumption 1, for all $t\in[0,1]$, the predictable quadratic variation process, as $n\to\infty$,
+$V_{n}(t):=\frac{1}{n}\sum_{k=0}^{\lfloor nt\rfloor-1}\mathbb{E}[(\Delta M_{i}^{(n)})^{2}|\mathcal{F}^{(n)}_{i-1}]=\frac{1}{n}\sum_{k=0}^{\lfloor nt\rfloor-1}U^{2}(\lambda_{i}^{(n)})\to\gamma^{2}t\ \text{in probability}.$
+When this condition is satisfied, the Lindeberg condition is equivalent to proving that $\max_{0\leq k\leq\lfloor nt\rfloor-1}\frac{1}{\sqrt{n}}|\Delta M_{i}^{(n)}|\to 0$ in probability.
+For any $\varepsilon>0$, applying the Markov inequality and Assumption 3 yields
+$\begin{split}\mathbb{P}(\max_{0\leq k\leq\lfloor nt\rfloor-1}\frac{1}{\sqrt{n}}|\Delta M_{i}^{(n)}|>\varepsilon)\leq\frac{1}{n^{2}\varepsilon^{4}}\sum_{k=0}^{\lfloor nt\rfloor-1}\mathbb{E}|\Delta M_{i}^{(n)}|^{4}\leq C_{\varepsilon}\frac{\lfloor nt\rfloor}{n^{2}}\to 0.\end{split}$
+The validity of these conditions implies that $S_{n}\Rightarrow\gamma W$ in $D([0,1])$. Consequently, the finite-dimensional distributions of $S_{n}$ converge to those of $\gamma W$.
+Step 2: Tightness in $\mathcal{C}^{\lambda}([0,1])$.
+According to Theorem 6 in ravckauskas2007holderian, tightness in the Hölder space requires two conditions: (i) tightness of the sequence of random variables $\{S_{n}(t)\}$ for each dyadic $t\in[0,1]$, and (ii) the uniform decay of the dyadic projections.
+The convergence of finite-dimensional distributions established in Step 1 ensures that the first condition holds.
+Therefore, it suffices to verify the second condition regarding the Schauder coefficients: for any $\varepsilon>0$,
+$\lim_{J\to\infty}\limsup_{n\to\infty}\mathbb{P}\Big(\sup_{j\geq J}2^{\lambda j}\max_{1\leq k\leq 2^{j}}|\delta_{j,k}(S_{n})|>\varepsilon\Big)=0,$ (A.2)
+where $\delta_{j,k}(S_{n})$ denotes the Schauder coefficients of $S_{n}$, defined by
+$\begin{split}\delta_{j,k}(S_{n})&=S_{n}(k/2^{j})-\frac{S_{n}((k-1)/2^{j})+S_{n}((k+1)/2^{j})}{2}\\ &=-\frac{1}{2}\left(\Delta S_{n}(k/2^{j})-\Delta S_{n}((k-1)/2^{j})\right),\end{split}$
+with $\Delta S_{n}(t):=S_{n}(t+2^{-j})-S_{n}(t)$. Elementary bounds yield
+$\left|\delta_{j,k}(S_{n})\right|\leq\left|\Delta S_{n}\left(\frac{k}{2^{j}}\right)\right|+\left|\Delta S_{n}\left(\frac{k-1}{2^{j}}\right)\right|\leq n^{-1/2}\sum_{i=\lfloor(k-1)n/2^{j}\rfloor}^{\lfloor(k+1)n/2^{j}\rfloor}\left|\Delta M_{i}^{(n)}\right|=:\Delta_{n,j,k}.$
+When $n<2^{j}$, sharper bounds apply: if the interval $(k-1)n/2^{j},(k+1)n/2^{j}]$ contains no integer points, $\delta_{j,k}(S_{n})=0$; otherwise,
+$\left|\delta_{j,k}(S_{n})\right|\leq\frac{n^{1/2}}{2^{j-1}}\left(\left|\Delta M_{\lfloor kn/2^{j}\rfloor}^{(n)}\right|\vee\left|\Delta M_{\lfloor(k-1)n/2^{j}\rfloor}^{(n)}\right|\right)=:\Delta_{n,j,k}^{*}$ (A.3)
+and at most $2n$ such intervals contain integer points.
+We split the probability estimate in (A.2) at the critical scale $j\approx\log_{2}n$. For $p>2$,
+$\begin{split}&\mathbb{P}\left(\sup_{j\geq J}2^{\lambda j}\max_{1\leq k\leq 2^{j}}|\delta_{j,k}(S_{n})|>\varepsilon\right)\\ &\leq\sum_{j\geq J}\sum_{1\leq k\leq 2^{j}}\mathbb{P}\left(|\delta_{j,k}(S_{n})|>\frac{\varepsilon}{2^{\lambda j}}\right)\\ &\leq\sum_{J\leq j\leq\lfloor\log_{2}n\rfloor}\sum_{1\leq k\leq 2^{j}}\frac{\mathbb{E}|\Delta_{n,j,k}|^{p}}{(\varepsilon 2^{-\lambda j})^{p}}+2n\sum_{j>\lfloor\log_{2}n\rfloor}\frac{\mathbb{E}|\Delta_{n,j,k}^{*}|^{p}}{(\varepsilon 2^{-\lambda j})^{p}}.\end{split}$ (A.4)
+Thus the key estimate relies on controlling the $p$-moments of $\Delta_{n,j,k}$ and $\Delta_{n,j,k}^{*}$. Let $I_{n,j,k}=[\lfloor(k-1)n/2^{j}\rfloor,\lfloor(k+1)n/2^{j}\rfloor]\cap\mathbb{Z}$.
+For the first part, by Rosenthal’s inequality, there exists a constant $C$ such that
+[…]
+Since $|I_{n,j,k}|=\lfloor(k+1)n/2^{j}\rfloor-\lfloor(k-1)n/2^{j}\rfloor+1\leq\frac{2n}{2^{j}}+2$, when $j\leq\lfloor\log_{2}n\rfloor$, we have
+$\mathbb{E}\left|\Delta_{n,j,k}\right|^{p}\leq C\left(n^{-1}|I_{n,j,k}|\right)^{p/2}\leq 2^{p}C2^{-jp/2}.$
+Consequently, the first summation in (A.4) satisfies
+$\sum_{J\leq j\leq\lfloor\log_{2}n\rfloor}\sum_{1\leq k\leq 2^{j}}\frac{\mathbb{E}\left|\Delta_{n,j,k}\right|^{p}}{(\varepsilon 2^{-\lambda j})^{p}}\leq\frac{2^{p}C}{\varepsilon^{p}}\sum_{J\leq j\leq\lfloor\log_{2}n\rfloor}2^{(1+\lambda p-p/2)j}.$
+For the second part, using (A.3), we have
+$2n\sum_{j>\lfloor\log_{2}n\rfloor}\frac{\mathbb{E}|\Delta_{n,j,k}^{*}|^{p}}{(\varepsilon 2^{-\lambda j})^{p}}\leq\frac{4C_{p}}{\varepsilon^{p}}n^{1+p/2}\sum_{j>\lfloor\log_{2}n\rfloor}2^{(\lambda-1)pj}.$
+Since $\lambda<1$, the geometric series $\sum_{j>\lfloor\log_{2}n\rfloor}2^{(\lambda-1)pj}$ converges and is of order $\mathcal{O}(n^{(\lambda-1)p})$.
+Therefore, the second summation is $\mathcal{O}(n^{1+\lambda p-p/2})$, which vanishes as $n\to\infty$, provided $\lambda<\frac{p-2}{2p}$.
+Combining these estimates yields
+$\limsup_{n\to\infty}\mathbb{P}\left(\sup_{j\geq J}2^{\lambda j}\max_{1\leq k\leq 2^{j}}\left|\delta_{j,k}(S_{n})\right|>\varepsilon\right)\leq\frac{2^{p}C}{\varepsilon^{p}}\frac{2^{(1+\lambda p-p/2)J}}{1-2^{1+\lambda p-p/2}}\to 0,\quad\text{as}\ J\to\infty.$
+Since $p>2$ can be taken arbitrarily large, tightness holds for all $\lambda<1/2$.
+A.5 Proof of Proposition 9
+Comparing (2.5) with (3.6), we have
+$\begin{split}&\left|\Lambda_{t}^{(n)}-\widetilde{\Lambda}_{t}^{(n)}\right|\\ &\leq\left|(1-a_{n})\theta\right|+\theta\left|\sum_{i=1}^{\lfloor nt\rfloor}(1-a_{n})\psi_{i}^{(n)}-\int_{0}^{t}f^{\alpha,\kappa}(u){\rm d}u\right|+\frac{\theta}{\omega_{n}}\left|\Delta M_{\lfloor nt\rfloor}^{(n)}\int_{t_{\lfloor nt\rfloor}}^{t}f^{\alpha,\kappa}(t-s){\rm d}s\right|\\ &+\frac{\theta}{n\omega_{n}}\left|\sum_{i=0}^{\lfloor nt\rfloor-1}\left[n(1-a_{n})\psi_{\lfloor nt\rfloor-i}^{(n)}-n\int_{t_{i}}^{t_{i+1}}f^{\alpha,\kappa}(t-s){\rm d}s\right]\Delta M_{i}^{(n)}\right|=I_{1}+I_{2}+I_{3}+I_{4}.\end{split}$
+It is evident that $I_{1}\to 0$.
+By (3.8) and Dini’s theorem, $I_{2}$ converges to zero uniformly.
+As for $I_{3}$, this term handles the boundary effect of the discretization interval.
+We first estimate the integral,
+$\left|\int_{t_{\lfloor nt\rfloor}}^{t}f^{\alpha,\kappa}(t-s){\rm d}s\right|\leq\int_{0}^{1/n}Cs^{\alpha-1}{\rm d}s=\frac{C}{\alpha}\left(\frac{1}{n}\right)^{\alpha}.$
+Thus,
+$\mathbb{E}|\sup_{t\in[0,1]}I_{3}|\leq\frac{\theta}{\omega_{n}}\mathbb{E}[\max_{k\leq k_{n}}|\Delta M_{k}^{(n)}|]\cdot\frac{C}{\alpha}n^{-\alpha}\lesssim n^{1/2-\alpha}\to 0.$
+Finally, we focus on $I_{4}$, the most critical component, which captures the discrepancy between the stepwise kernel and the average density of $f^{\alpha,\kappa}$ over small intervals.
+Noting that $n(1-a_{n})\psi_{\lfloor nt\rfloor}^{(n)}/\rho_{n}(t)=a_{n}\to 1$, we will henceforth replace $n(1-a_{n})\psi_{\lfloor nt\rfloor-i}^{(n)}$ with $\rho_{n}(t-t_{i})$.
+Using a right-endpoint approximation, we decompose
+[…]
+By Proposition 4, $\|K_{n}\|_{2}\to 0$, and moreover $\|K_{n}\|_{2}\lesssim n^{1/2-\alpha}$ according to wang2025rough.
+Since $f^{\alpha,\kappa}(t)\sim t^{\alpha-1}$ near zero, standard regularity arguments imply that $\|D_{n}\|_{2}\to 0$, with the same rate $n^{1/2-\alpha}$.
+Hence Minkowski’s inequality gives
+$\|G_{n}\|_{2}\leq\|K_{n}\|_{2}+\|D_{n}\|_{2}\lesssim n^{1/2-\alpha}.$
+Denote $C_{n}=\frac{\theta}{n\omega_{n}}\sim n^{-1/2}$ and $G_{i,n}=G(t_{i})$. We define the empirical $L^{2}$-norm of $(G_{n,i})_{i\geq 0}$ is actually the discrete version of $\|G_{n}\|_{2}$:
+$\|G\|_{n,2}:=\left(\frac{1}{n}\sum_{i=0}^{n-1}G_{i,n}^{2}\right)^{1/2}\sim\|G_{n}\|_{2}\lesssim n^{1/2-\alpha}.$
+Thus, we can represent $I_{4}$ through
+$I_{4}=C_{n}\left|\sum_{i=1}^{\lfloor nt\rfloor}G_{n,i}\Delta M_{\lfloor nt\rfloor-i}^{(n)}\right|=:|\mathfrak{G}_{t}^{(n)}|.$
+The process $\mathfrak{G}^{(n)}$ is a weighted sum of martingale difference but not not itself a martingale, so classical maximal inequalities do not apply directly.
+Inspired by the factorization method for stochastic convolutions introduced in da2014stochastic, we employ a discrete counterpart leveraging the semigroup property of fractional difference operators.
+Let $\delta\in(1/2,\alpha)$ be a fixed parameter.
+We explicitly define the smoothing sequence $a=(a_{k})_{k\geq 0}$ via the generalized binomial coefficients corresponding to the fractional order $\alpha-\delta$:
+$a_{k}:=(-1)^{k}\binom{-(\alpha-\delta)}{k}=\frac{\Gamma(k+\alpha-\delta)}{\Gamma(k+1)\Gamma(\alpha-\delta)}\sim\frac{1}{\Gamma(\alpha-\delta)}k^{\alpha-\delta-1},\quad\text{as }k\to\infty.$
+Given the error kernel $G$, we define the sequence $b=(b_{k,n})_{k\geq 1}$ as the unique solution to the discrete convolution equation G=abG=a*b.
+The existence and uniqueness of $b$ are guaranteed by the invertibility of the lower-triangular Toeplitz matrix since $a_{0}=1$.
+To control the magnitude of $b$, consider $c$ as the convolution inverse of $a$, with generating function
+$C(z)=\frac{1}{A(z)}=\frac{1}{(1-z)^{-(\alpha-\delta)}}=(1-z)^{\alpha-\delta},$
+so that $c_{k}=\binom{\alpha-\delta}{k}\sim k^{-(\alpha-\delta)-1}$. Since $\alpha>\delta$, which implies $c\in l^{1}$. By the discrete Young’s inequality, the empirical $L^{2}$-norm of $b$ satisfies:
+$\|b\|_{n,2}:=\left(\frac{1}{n}\sum_{i=1}^{n}b_{i,n}^{2}\right)^{1/2}\leq\|c\|_{l^{1}}\|G\|_{n,2}\lesssim n^{1/2-\alpha}.$
+Using the factorization $G_{k,n}=\sum_{j=1}^{k}a_{k-j}b_{j,n}$ and exchanging the order of summation, we obtain:
+$\mathfrak{G}_{t}^{(n)}=C_{n}\sum_{j=1}^{\lfloor nt\rfloor}\left(\sum_{i=1}^{j}a_{j-i}b_{i,n}\right)\Delta M_{\lfloor nt\rfloor-j}^{(n)}=\sum_{j=0}^{\lfloor nt\rfloor-1}a_{j}\mathcal{Y}_{\lfloor nt\rfloor-j},$
+where the auxiliary process $\mathcal{Y}$ is defined as $\mathcal{Y}_{k}:=\sum_{i=1}^{k}b_{i,n}C_{n}\Delta M_{k-i}^{(n)}$. Applying Hölder’s inequality with $\frac{1}{p}+\frac{1}{q}=1$, we have
+$\sup_{t\in[0,1]}|\mathfrak{G}_{t}^{(n)}|\leq\left(\sum_{k=0}^{n-1}|a_{k}|^{q}\right)^{1/q}\cdot\left(\sum_{k=0}^{n-1}|\mathcal{Y}_{k}|^{p}\right)^{1/p}.$
+We choose $q>\frac{1}{1-(\alpha-\delta)}$, ensuring $\left(\sum_{k=0}^{n-1}|a_{k}|^{q}\right)^{1/q}\leq\|a\|_{l^{q}}<\infty$.
+It remains to analyze the term involving $\mathcal{Y}$, which is still not a martingale.
+But for a fixed $k$, $\mathcal{Y}_{k}$ can be regarded as the terminal value of a discrete martingale with respect to the index $i$.
+Using the Burkholder-Davis-Gundy (BDG) inequality, we obtain
+$\mathbb{E}|\mathcal{Y}_{k}|^{p}\leq C_{p}\mathbb{E}\left(\sum_{i=1}^{k}|b_{i,n}|^{2}C_{n}^{2}|\Delta M_{k-i}^{(n)}|^{2}\right)^{p/2}.$
+Noting that $C_{n}^{2}\sim n^{-1}$, by Minkowski’s inequality and the Assumption 3, we deduce
+$\left\|\sum_{i=1}^{k}|b_{i,n}|^{2}C_{n}^{2}|\Delta M_{k-i}^{(n)}|^{2}\right\|_{L^{p/2}}\leq\sum_{i=1}^{k}|b_{i,n}|^{2}C_{n}^{2}\|\Delta M_{k-i}^{(n)}\|_{L^{p}}^{2}\lesssim\frac{1}{n}\sum_{i=1}^{k}|b_{i,n}|^{2}.$
+Consequently, utilizing the bound $\|b\|_{n,2}\lesssim n^{1/2-\alpha}$, we estimate the total $p$-th moment:
+$\mathbb{E}\sum_{k=0}^{n-1}|\mathcal{Y}_{k}|^{p}\lesssim\sum_{k=0}^{n-1}\left(\frac{1}{n}\sum_{i=1}^{k}|b_{i,n}|^{2}\right)^{p/2}\leq n\left(\|b\|_{n,2}^{2}\right)^{p/2}\lesssim n\cdot(n^{1/2-\alpha})^{p}=n^{1+p(1/2-\alpha)}.$
+For convergence, we require the exponent $1+p(1/2-\alpha)<0$, which implies $p>\frac{1}{\alpha-1/2}$.
+On the other hand, the constraint for $q>\frac{1}{1-(\alpha-\delta)}$ implies $p<\frac{1}{\alpha-\delta}$.
+Since $\delta>1/2$, a feasible $p$ exists.
+For such $p$, $\mathbb{E}\sum_{k=0}^{n-1}|\mathcal{Y}_{k}|^{p}\to 0$, which implies $\left(\sum_{k=0}^{n-1}|\mathcal{Y}_{k}|^{p}\right)^{1/p}\to 0$ in probability.
+Finally, we conclude
+$\sup_{t\in[0,1]}|\mathfrak{G}_{t}^{(n)}|\leq\|a\|_{l^{q}}\cdot\left(\sum_{k=0}^{n-1}|\mathcal{Y}_{k}|^{p}\right)^{1/p}\to 0\quad\text{in probability}.$
+In summary, $\Lambda_{t}^{(n)}$ and $\widetilde{\Lambda}_{t}^{(n)}$ are asymptotically indistinguishable.
+Appendix B The proof in Section 4
+B.1 Proof of Theorem 10
+The proof necessitates establishing the joint weak convergence of the return process and the volatility process.
+We firstly verify the Theorem 1 so that the $\Lambda^{(n)}$ converges.
+Denote $\mathbb{E}_{k-1}[\cdot]:=\mathbb{E}[\cdot|\mathcal{F}_{k-1}]$.
+By the properties of GED, for $p\geq 2$, the moments of the innovation term $\varepsilon_{k}$ satisfy:
+$\mathbb{E}_{k-1}[|\varepsilon_{k}|^{p}]=\frac{2^{p/\nu}}{\nu\Gamma(1+1/\nu)}\Gamma\left(\frac{p+1}{\nu}\right)<\infty.$
+Recall that the score function is defined as $\nabla_{k}=\frac{\nu}{2}|\varepsilon_{k}|^{\nu}-1$. We have,
+$\mathbb{E}[|\nabla_{k}|^{p}]\leq 2^{p-1}\left[\left(\frac{\nu}{2}\right)^{p}\mathbb{E}|\varepsilon_{k}|^{\nu p}+1\right]<\infty.$
+In particular, the second moment is given by
+$U^{2}(\lambda_{k})=\mathbb{E}_{k-1}[\nabla_{k}^{2}]=\frac{\nu^{2}}{4}\mathbb{E}_{k-1}[|\varepsilon|^{2\nu}]-\nu\mathbb{E}_{k-1}[|\varepsilon|^{\nu}]+1=\nu.$
+The dynamic of $\Lambda_{t}$ then follows from Theorem 1.
+Then we establish thge joint convergence of driving Noises. Consider the vector of partial sums representing the driving noises for the returns and the volatility parameter:
+$\mathbf{M}^{(n)}_{t}=\left(B^{(n)}_{t},W^{(n)}_{t}\right):=\left(\sum_{i=1}^{\lfloor nt\rfloor}\frac{\varepsilon_{i}}{\sqrt{\chi n}},\sum_{i=1}^{\lfloor nt\rfloor}\frac{\nabla_{i}}{\sqrt{\nu n}}\right).$
+Since the sequence $\{\varepsilon_{i}\}$ is i.i.d., we apply the multidimensional Donsker’s invariance principle.
+Note that $\mathbb{E}_{n-1}(\varepsilon_{n})=0$ and by the properties of score implies $\mathbb{E}_{n-1}(\nabla_{n})=0$ and $\mathbb{E}_{n-1}(\varepsilon_{n}\nabla_{n})=0$.
+Consequently, $\mathbf{M}^{(n)}$ converges weakly in the Skorohod space $\mathcal{D}([0,1],\mathbb{R}^{2})$ to a two-dimensional Brownian motion $(B,W)$ with an identity covariance matrix.
+Furthermore, since $\Lambda^{(n)}$ is adapted to the filtration generated by $\mathbf{M}^{(n)}$, the joint convergence holds:
+$(\Lambda^{(n)},B^{(n)})\Rightarrow(\Lambda,B)\quad\text{in }\mathcal{D}([0,1],\mathbb{R}^{2}).$ (B.1)
+Finally, we address the convergence of $X^{(n)}$.
+By the continuous mapping theorem, since $\Lambda^{(n)}\Rightarrow\Lambda$, we have $H^{(n)}={\rm e}^{\Lambda^{(n)}}\Rightarrow e^{\Lambda}=:H$.
+We now invoke the convergence theorem for stochastic integrals (see, , e.g., kurtz1991weak).
+The conditions required are: (i) The joint convergence of the integrand and integrator, which is already satisfied in (B.1); and (ii) The uniform tightness of integrators, which is equivalent to verify the quadratic variation is stochastically bounded.
+We observe that
+$\mathbb{E}([B^{(n)},B^{(n)}]_{t})=\frac{1}{n}\sum_{i=1}^{\lfloor nt\rfloor}\mathbb{E}(\varepsilon_{i}^{2}/\chi)\leq t.$
+Therefore, by Markov’s inequality, for any $K>0$,
+$\mathbb{P}([B^{(n)},B^{(n)}]_{t}>K)\leq\frac{\mathbb{E}([B^{(n)},B^{(n)}]_{t})}{K}\leq\frac{t}{K}\to 0,\quad\text{as}\ K\to\infty.$
+This confirms the uniform tightness. Thus, the sequence of stochastic integrals converges weakly:
+$X^{(n)}_{t}=\int_{0}^{t}H^{(n)}_{s-}{\rm d}B^{(n)}_{s}\Rightarrow\int_{0}^{t}{\rm e}^{\Lambda_{s}}{\rm d}B_{s}=X_{t}$
+in the Skorohod topology. Combining this with the marginal convergence of $\Lambda$, we conclude that the pair $(X^{(n)},\Lambda^{(n)})$ converges weakly to $(X,\Lambda)$ as stated.